### v2

# An Efficient Classical Algorithm for Simulating Short Time 2D Quantum Dynamics

Abstract
Efficient classical simulation of the Schrödinger equation is central to quantum mechanics, as it is crucial for exploring complex natural phenomena and understanding the fundamental distinctions between classical and quantum computation. Although simulating general quantum dynamics is BQP-complete, tensor networks allow efficient simulation of short-time evolution in 1D systems. However, extending these methods to higher dimensions becomes significantly challenging when the area law is violated. In this work, we tackle this challenge by introducing an efficient classical algorithm for simulating short-time dynamics in 2D quantum systems, utilizing cluster expansion and shallow quantum circuit simulation. Our algorithm has wide-ranging applications, including an efficient dequantization method for estimating quantum eigenvalues and eigenstates, simulating superconducting quantum computers, dequantizing quantum variational algorithms, and simulating constant-gap adiabatic quantum evolution. Our results reveal the inherent simplicity in the complexity of short-time 2D quantum dynamics and highlight the limitations of noisy intermediate-scale quantum hardware, particularly those confined to 2D topological structures. This work advances our understanding of the boundary between classical and quantum computation and the criteria for achieving quantum advantage.
I Introduction
A longstanding open question is understanding the classical computational complexity of simulating quantum dynamics. While simulating general quantum dynamics is BQP-complete and therefore unlikely to be efficiently classically simulated, certain quantum systems with specific geometric structures allow for efficient classical algorithms. For example, short-time dynamics of (quasi-)1D systems that satisfy the area law can be efficiently simulated using matrix product states (MPS) [1, 2, 3, 4]. However, extending these methods to 2D systems is more challenging, and whether 2D systems generally satisfy the area law remains an open question [5]. Although 2D frustration-free gapped Hamiltonians have been shown to obey the area law [6], more complex 2D systems pose difficulties for MPS-based approaches. Efficiently representing such states using MPS could imply a collapse of the polynomial hierarchy [7]. Theoretical results [8, 9, 10] show that certain 2D systems satisfying the area law are quantumly hard to simulate. Furthermore, it has been demonstrated that preparing 2D area law states can be PostBQP-hard [8, 11], where PostBQP machines, which allow postselection, are believed to be more powerful than QMA. These PostBQP-hard results also apply to the preparation of projected entangled pair states (PEPS), which extend MPS to higher dimensions and provide a framework for describing quantum states in 2D systems. Classical numerical simulations face difficulties with PEPS algorithms, especially those based on variational principles. The tensor contraction process in PEPS is classically hard [10, 12], and normalizing PEPS has been proven to be an undecidable problem [13], complicating practical optimization. Most of the above results focus on the preparation of 2D quantum states. An ostensibly simpler task is computing the expectation value of observables on such states. However, even this problem is #P-hard for Hermitian observables when considering relative error approximations [14]. As a result, the efficient classical simulability of 2D quantum dynamics remains a significant open problem.
In modern physics, 2D Hamiltonians are fundamental in both theoretical research and practical applications. They are essential for studying quantum systems in two-dimensional spaces, offering critical insights into properties of quantum states and entanglement [15, 16]. They also serve as key models for phenomena such as the quantum and anomalous Hall effects [17, 18], capturing the behavior of 2D materials under magnetic fields. In condensed matter physics, the 2D Fermi-Hubbard model plays a central role in understanding complex phenomena like superconductivity and magnetism [19, 20, 21], aiding the development of new superconducting materials and magnetic storage technologies. 2D Hamiltonians are also widely applied in materials science, particularly in characterizing the electronic and optical properties of two-dimensional materials like graphene [22]. By studying the ground states of 2D Hamiltonians, researchers can predict material behavior and design functional materials for advanced technologies such as electronics and sensors [23]. In quantum computing, 2D quantum systems provide a foundation for scalable superconducting quantum processors [24, 25]. These processors have already demonstrated applications in random circuit sampling [26, 24, 25], quantum many-body system dynamics [27, 28, 29], entanglement witnessing [30], and quantum chemistry problems [31, 32, 33, 34].
Given the importance of 2D Hamiltonians in both theoretical research and practical applications, efficiently simulating the dynamics of 2D quantum systems and predicting their ground state properties is a crucial challenge. In this paper, we address this by presenting an efficient classical algorithm for simulating 2D quantum dynamics. Consider a set of $n$-qubit 2D local Hamiltonians $\{H^{(k)}\}_{k=1}^{K}$ and a time series $\vec{t}=\{t_{k}\}_{k=1}^{K}$, along with an $n$-qubit Hermitian observable $O=O_{1}\otimes\cdots\otimes O_{n}$. The mean value of $O$ with respect to the state $|\psi(\vec{t})\rangle=e^{-iH^{(1)}t_{1}}\cdots e^{-iH^{(K)}t_{K}}|0^{n}\rangle$ is $\mu(\vec{t})=\langle\psi(\vec{t})|O|\psi(\vec{t})\rangle$. When $K$ and $t=\max\{t_{k}\}_{k=1}^{K}$ are constants, our classical algorithm provides a $\epsilon$-approximation to $\mu(\vec{t})$ with a running time of $\mathcal{O}(n^{\log(n/\epsilon)})$, which is significantly closer to polynomial than exponential complexity. Thus, for medium-scale systems (e.g., $n\leq 10^{3}$), our algorithm can efficiently solve the quantum mean value problem for 2D quantum systems. This approach non-trivially extends prior works, including Ref. [35] from local to general global observables and Ref. [36] from discrete circuits to continuous Hamiltonian dynamics.
Our results have several significant applications. First, for 2D Hamiltonians exhibiting certain symmetries, such as particle or spin conservation, our algorithm can solve the guided local Hamiltonian problem without the bounded Hamiltonian norm assumption used in Ref. [37]. Notably, our classical algorithm provides only constant-precision ground state energy estimates, thus remaining consistent with the BQP-complete result of Ref. [38]. Furthermore, since current superconducting quantum computing platforms employ a 2D topological structure for qubit connections [24, 25], our algorithm is naturally suited for simulating superconducting quantum computations with constant evolution times. Interestingly, for non-constant evolution times, recent studies have shown that noisy quantum circuits with constant noise rates are classically simulable [39, 40]. These findings are crucial in identifying the conditions for achieving quantum advantage on superconducting hardware. Moreover, in the NISQ era, variational algorithms such as the variational quantum eigensolver (VQE) and quantum approximate optimization algorithm (QAOA) [41, 42, 43, 44, 45, 46] have been widely used to approximate ground states of quantum systems using variational principles. Our classical algorithm efficiently simulates constant-layer VQE and QAOA driven by 2D Hamiltonians. Additionally, we show that it can simulate short-time adiabatic quantum computation under constant evolution time. These results suggest that, for 2D quantum systems with constant evolution times, current superconducting quantum computers may not offer an exponential quantum advantage. Instead, classical algorithms can simulate various behaviors of 2D quantum systems in quasi-polynomial time. This work deepens the understanding of the provable advantages of quantum computing and refines the boundary between classical and quantum computational power.
Figure 1: (a) A quantum circuit representation on the quantum dynamics mean value. The quantum dynamics is governed by 2D Hamiltonians $\{H^{(k)}\}_{k=1}^{K}$ and corresponding time $\{t_{k}\}_{k=1}^{K}$, which acts on the initial state $|0^{n}\rangle$. (b) Visualization on approximating $U_{i}(\vec{t})O_{i}U^{\dagger}_{i}(\vec{t})$ by using the cluster expansion method. In this visualization, each grey point represents a single qubit and the red dot circle represents a Hermitian term $h_{X}$. Without loss of generality, we assume that $h_{X}$ is $2$-local in the visualization. The approximation is essentially a linear combination of ${\rm poly}(n)$ matrices induced by connected clusters, and it is applied to $M\leq\mathcal{O}(e^{K\mathfrak{d}t}\log(2n/\epsilon))$ qubits. (c) The analytic continuation method provides a paradigm to approximate $U_{i}(\vec{t})O_{i}U^{\dagger}_{i}(\vec{t})$ for general $\max\{t_{k}\}\leq\mathcal{O}(1)$. We leave details to Lemma 5. (d) Visualization of partitions of 2D grid into regions $R_{1}$ and $R_{2}$ for $M=1$, where $M$ represents the short-side length of each sub-region $R_{i}(l)$ ($i\in\{1,2\})$) highlighted by dark or light blue points. Their supports ${\rm supp}(V_{R_{i}}(l))$ are depicted by $(\sqrt{n}\times 4M)$-size black dot circles, and it is shown that ${\rm supp}(V_{R_{i}}(l))\cap{\rm supp}(V_{R_{i}}(l^{\prime}))=\emptyset$ for all $l\neq l^{\prime}$. (e) The 2D dynamics mean value problem 1 can be employed to predict 2D ground state energy (such as the Fermi-Hubbard model), simulate superconducting quantum computation results, the outputs of VQE and QAOA, and short-time adiabatic quantum computation.
II Problem Statement
We consider $n$-qubit $2$D geometrical local Hamiltonian
$H=\sum\limits_{X\in S}\lambda_{X}h_{X},$ (1)
where $S$ represents a set of subsystems, real-valued coefficient $\left|\lambda_{X}\right|\leq 1$ and $h_{X}$ represents a Hermitian operator non-trivially acting on the 2D geometrical local qubit set $X\subset S$. Without loss of generality, we assume the operator norm of each $h_{X}$ satisfies $\|h_{X}\|\leq 1$. Note that the norm of the total Hamiltonian $\|H\|$ is not bounded by 1.
To characterize the locality and correlations presented by the Hamiltonian, we introduce the associated interaction graph $G$ to depict overlaps of operators contained in $H$ [47, 35]. Specifically, given the Hamiltonian terms $\{h_{X}\}_{X\subset S}$, the interaction graph $G$ is a simple graph with vertex set $\{h_{X}\}_{X\subset S}$. An edge exists between $h_{X}$ and $h_{X^{\prime}}$ if $X\cap X^{\prime}\neq\emptyset$, and we denote the degree $\mathfrak{d}(h_{X})$ of a vertex $h_{X}$, which is the number of edges incident to it. The maximum degree among all vertexes within the interaction graph $G$ is denoted by $\mathfrak{d}=\max_{h_{X}\in H}\left\{\mathfrak{d}(h_{X})\right\}$. Further details on the definitions of the 2D Hamiltonian and the interaction graph are provided in Appendix B.
In this article, we consider the computation of expectation values at the output of a $K$-step Hamiltonian dynamics.
Problem 1 ($K$-step Quantum Dynamics Mean Value).
Consider $K$ local Hamiltonians $\{H^{(1)},H^{(2)},\cdots,H^{(K)}\}$ defined on a 2D plane, and a global observable $O=O_{1}\otimes\cdots\otimes O_{n}$ with the operator norm $\|O_{i}\|\leq 1$ for $i\in[n]$. The $K$-step quantum mean value is defined by
$\mu(\vec{t})=\langle 0^{n}|\left(\prod\limits_{k=1}^{K}e^{-iH^{(k)}t_{k}}\right)^{\dagger}O\left(\prod\limits_{k=1}^{K}e^{-iH^{(k)}t_{k}}\right)|0^{n}\rangle,$ (2)
where evolution time series $\vec{t}=\{t_{1},\cdots,t_{K}\}$. The target is to provide an estimation $\hat{\mu}(\vec{t})$ such that $\left|\mu(\vec{t})-\hat{\mu}(\vec{t})\right|\leq\epsilon$.
We note that many problems can be transformed into solving the Problem 1. For example, the 2D Hamiltonian ground state problem is one of the most significant problems in condensed matter physics, which plays a vital role in expressing superconductivity, magnetism, and other phenomena. Assuming ${\rm BPP}\neq{\rm BQP}$, estimating the ground state energy within $\delta\leq 1/{\rm poly}(n)$ additive error may be classically hard [38]. It is interesting to note that the local-guided Hamiltonian problem can be solved by using Problem 1 as a subroutine. Given an initial quantum state $|\psi_{c}\rangle$ with an overlap $p_{0}$ to the ground state, the fundamental idea is to reconstruct the cumulative distribution function
$C(x)=\sum_{\left|j\right|\leq l}\hat{F}_{j}e^{ijx}\langle\psi_{c}|e^{-ijH}|\psi_{c}\rangle$
associated with the Hamiltonian $H$ [48]. Here, the jump of $C(x)$ (from zero to nonzero) determines the ground state energy, $\hat{F}_{j}$ represents the Fourier transformation of the Heaviside function, and the maximum evolution time $l\leq\mathcal{O}(\delta^{-1}\log(p_{0}))$. As a result, samples from the Loschmidt echo $\langle\psi_{c}|e^{-ijH}|\psi_{c}\rangle$ suffice to provide a quantum ground state energy estimation. In general, classically simulate the Loschmidt echo $\langle\psi_{c}|e^{-ijH}|\psi_{c}\rangle$ is challenging for $j\in\mathcal{O}(1)$, however, when the Hamiltonian $H$ satisfies certain symmetries (such as particle number preserve), the Loschmidt echo can be equivalently achieved by solving a quantum dynamics mean value problem (Problem 1). We refer more detailed discussions to Sec. V.1 and Appendix. L.
Furthermore, Problem 1 can be used to simulate the dynamical behavior of both digital and analog superconducting quantum computers, where each quantum circuit layer can be equivalently achieved by a Hamiltonian dynamics driven by a 2D Hamiltonian (given by Eq. 1). When the observable $O=|x\rangle\langle x|$ for some $x\in\{0,1\}^{n}$, $\mu(\vec{t})$ may represent the probability of sample $x$ from the quantum state $\prod_{k=1}^{K}e^{iH^{(k)}t_{k}}|0^{n}\rangle$. When $O=\otimes_{i=1}^{n}P_{i}$ ($P_{j}\in\{I,X,Y,Z\}$), it enables us to obtain any linear property of the quantum system.
The variational quantum algorithms (VQE) [41, 49] are widely studied and experimentally verified on superconducting quantum platforms [31, 33]. VQE aims at finding the ground state energy of quantum many-body systems by minimizing the energy function $E_{U}=\langle 0^{n}|U^{\dagger}H_{e}U|0^{n}\rangle,$ where $H_{e}$ represent the quantum lattice model or electronic Hamiltonian, and $U|0^{n}\rangle$ can represent the parameterized quantum trial state prepared by the Hamiltonian-Variational (HV) ansatz [50]. Using the superfast encoding method [51, 52, 53], $2$D local fermionic Hamiltonians (such as $2$D Fermi-Hubbard model) can be encoded by a sum of local Pauli operators [34], and the energy $E_{U}$ is thus a sum of mean values of the form given by Problem 1. Similarly, we study instances of the QAOA problem on constant-regular graphs [42, 54, 55, 56]. The MaxCut is a prototypical discrete optimization issue characterized by a low and fixed node degree along with a high dimension. It cannot be simply mapped to a planar architecture and is more closely related to problems of industrial value. When $G_{\rm QAOA}(V,E)$ represents a regular graph with constant vertex degree, the MaxCut problem aims at dividing all vertices into two disjoint sets such that maximize the number of edges that connect the two sets. The QAOA algorithm encodes the solution to the ground state of a local Hamiltonian $H_{A}=-\sum_{(i,j)\in E}\frac{1}{2}(\mathbb{I}-Z_{i}Z_{j})$ whose corresponding interaction graph has the constant degree. Then the HV ansatz $U$ is used to approximate the solution via minimizing the energy function $E_{U}=\langle+^{n}|U^{\dagger}H_{A}U|+^{n}\rangle$.
When considering time-dependent quantum dynamics, the adiabatic quantum process plays an important role in quantum physics ranging from quantum state preparation [57, 58], optimization [59, 60, 61], and many-body physics [62, 63]. For short-time adiabatic dynamics, these studies are closely related to the definition of quantum phases [62] and thus can be extended to the numerical study of properties of certain quantum phases. Here, we consider the classical estimation of the expectation value of observable with the constant-time adiabatic quantum process: $\braket{O(t)}=\bra{\psi}U^{\dagger}(t)OU(t)\ket{\psi}$, where $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$ is the unitary that defines the adiabatic dynamics, $O$ is a $k$-local observable with $k$ a constant and |$\ket{\psi}$ is a product state. The setting holds physical relevance for states adiabatically connected to a product state that is in the symmetry-protected topological (SPT) order [64, 65]. Our simulation methods could thus be made to facilitate the study of properties regarding SPT and other short-range entangled systems.
Given the close relationship between Problem 1 and the significant computational problems, it is natural to question whether Problem 1 can be efficiently solved by using classical computers.
III Main Results
Our main result is to provide a quasi-polynomial classical algorithm for solving Problem 1.
Theorem 1.
Given $K$ Hamiltonians $\{H^{(1)},\cdots,H^{(K)}\}$ defined on a $2D$ plane with $n$ qubits, any observable $O=O_{1}\otimes\cdots\otimes O_{L}$ with $\|O_{i}\|\leq 1$ and locality $L\leq n$, and a time series $\vec{t}=(t_{1},\cdots,t_{K})$, there exists a classical algorithm that outputs an approximation $\hat{\mu}(\vec{t})$ such that $\left|\mu(\vec{t})-\hat{\mu}(\vec{t})\right|\leq\epsilon$ with a run time of at most
$\mathcal{O}\left(\frac{n}{\epsilon^{2}}\left(\frac{2L}{\epsilon}\right)^{e^{2\pi eK\mathfrak{d}t}\log(2L/\epsilon)}\right),$ (3)
where $t=\max\{t_{k}\}_{k=1}^{K}$ and $\mathfrak{d}$ represents the maximum degree of the related interaction graphs.
We leave proof details and relevant lemmas to Section D. Our finding gives evidence that when the locality is $L=n$ and evolution time satisfies $t\leq\mathcal{O}(\log\log(n))$, the mean value $\mu(\vec{t})$ can be approximated by a quasi-polynomial classical algorithm within $\epsilon$-additive error. The ‘quasi-polynomial algorithm’ represents programs that return results in $\mathcal{O}(2^{\log^{a}(n)})$ running time for some fixed constant $a$, which is a significant improvement over algorithms with exponential running time. Notably, our result uses different techniques compared to Ref. [37]. Therefore, our algorithm does not depend on the operator norm assumption $\|H\|\leq 1$ as adopted in Ref. [37], which is not satisfied even for the simplest Ising Hamiltonian.
Furthermore, we note that the Problem 1 may not be trivially solved by the method proposed by S. Bravyi et al. [36], which proposed a classical algorithm for simulating the quantum mean value problem for general classes of quantum observables $O=O_{1}\otimes\cdots\otimes O_{n}$ and $2$D constant-depth quantum circuits $U$. Specifically, they divide $UOU^{\dagger}=\prod_{i=1}^{n}UO_{i}U^{\dagger}$ into two operators $U_{A}U_{B}$, where $U_{A}$ and $U_{B}$ can be classically simulated easily based on the “casual principle” given by the lightcone. By utilizing the classical Monte Carlo method, they can efficiently simulate the value $\langle 0^{n}|U_{A}U_{B}|0^{n}\rangle$ which approximates the quantum mean value. However, when the unitary $U$ is given by a Hamiltonian dynamics $e^{-iHt}$, the fundamental challenge arises because $e^{-iHt}O_{i}e^{iHt}$ may not be easily computed by the casual principle which only considers quantum gates in the light-cone of $UO_{i}U^{\dagger}$. The reason stems from the relatively large quantum circuit depth taken by the Hamiltonian simulation method. For example, the general Trotter-Suzuki method would require ${\rm poly}(\|H\|,t^{O(1)},1/\epsilon^{O(1)})$ [66]-depth quantum circuit. When the Hamiltonian norm $\|H\|=\Theta(n)$, the corresponding deep quantum circuit $U$ will introduce long-range correlations, and computing $UO_{i}U^{\dagger}$ by casual principle would be classically hard. Even the most advanced quantum simulation algorithm for $2$D local Hamiltonians requires quantum circuit depth $\mathcal{O}(t\log^{3}(nt/\epsilon))$-depth quantum circuit [67], and it takes $\mathcal{O}(2^{t^{2}\log^{6}(nt/\epsilon)})$ computational complexity to compute $UO_{i}U^{\dagger}$.
In this article, we present a method capable of overcoming the above obstacle, by combining the cluster expansion method [47, 35] and analytic continuation. The proposed method can efficiently approximate $e^{-iHt}O_{i}e^{iHt}$ in polynomial time. Our method nontrivial extends and generalizes Ref. [36] in solving quantum mean value problems when $U$ is given by a series of Hamiltonian dynamics. We will outline the classical simulation algorithm in Sec. IV.
As an interesting byproduct, we find that the cluster expansion also provides a novel way to characterize the quantum circuit complexity lower bound on the Hamiltonian dynamics problem which linearly grows with parameters $K$ and $t$. The complexity of a quantum circuit $C(U)$ characterizes the necessary quantum circuit depth required to synthesize $U$ [68, 69], meanwhile, the complexity has deep connections to high-energy physics, quantum many-body systems, and black-hole physics [70, 71, 72, 73, 74]. Compared with the previous art [75], our result holds for a large class of Hamiltonians which does not depend on the translation invariant constraint. We leave proof details to the Appendix F.
Corollary 1.
Given $K$ Hamiltonians $\{H^{(1)},\cdots,H^{(K)}\}$ defined on a two-dimensional plane with $n$ qubits and a time series $\vec{t}=(t_{1},\cdots,t_{K})$, we denote $U(\vec{t})=\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}}$. Then the quantum circuit depth complexity of $U(\vec{t})$ satisfies
$C(U(\vec{t}))\geq\Omega\left(K\mathfrak{d}t\right),$ (4)
where the constant time $t=\max\{t_{k}\}_{k=1}^{K}$, $\mathfrak{d}$ represents the maximum degree of the related interaction graphs.
IV Classical Algorithm for 2D Quantum Dynamics Mean Value
Let us outline the proposed classical algorithm. Denote the time series $\vec{t}=(t_{1},\cdots,t_{K})$, a Hamiltonian dynamics operator $U(\vec{t})=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}}$ driven by Hamiltonians $\{H^{(1)},\cdots,H^{(K)}\}$ and the observable $O=O_{1}\otimes\cdots\otimes O_{n}$, the quantum dynamics mean value can be equivalently computed by
$\begin{split}\mu(\vec{t})&=\langle 0^{n}|U^{\dagger}(\vec{t})\left(O_{1}\otimes\dots\otimes O_{n}\right)U(\vec{t})|0^{n}\rangle\\ &=\langle 0^{n}|\left(U^{\dagger}(\vec{t})O_{1}U(\vec{t})\right)\left(U^{\dagger}(\vec{t})O_{2}U(\vec{t})\right)\cdots\left(U^{\dagger}(\vec{t})O_{n}U(\vec{t})\right)|0^{n}\rangle\\ &=\langle 0^{n}|U_{1}(\vec{t})U_{2}(\vec{t})\cdots U_{n}(\vec{t})|0^{n}\rangle,\end{split}$ (5)
where $U_{i}(\vec{t})=U^{\dagger}(\vec{t})O_{i}U(\vec{t})$. Our first step aims to approximate $U_{i}(\vec{t})$ by $V_{i}(\vec{t})$ such that $\|U_{i}(\vec{t})-V_{i}(\vec{t})\|\leq\mathcal{O}(\epsilon/2n)$ under the operator norm. Here, the operator $V_{i}(\vec{t})$ is essentially a linear combination of ${\rm poly}(n)$ matrices which nontrivial perform on at most $\mathcal{O}(e^{K\mathfrak{d}t}\log(2n/\epsilon))$ qubits, given by the cluster expansion method (as shown in Fig. 1. b). Meanwhile, Lemma 5 demonstrated that $V_{i}(\vec{t})$ can be efficiently computed by a classical algorithm, by using the analytic continuation tools (Fig. 1. c). We leave technical details to Eq. 38 and Appendix C. After approximating $U_{i}(\vec{t})$ by operator $V_{i}(\vec{t})$ for index $i\in[n]$, the mean value $\mu(\vec{t})$ can be approximated by $\hat{\mu}(\vec{t})=\langle 0^{n}|V_{1}(\vec{t})\cdots V_{n}(\vec{t})|0^{n}\rangle$ such that $\left|\mu(\vec{t})-\hat{\mu}(\vec{t})\right|\leq\epsilon/2$.
The second step applies the causality principle and the lightcone of $V_{i}(\vec{t})$ to assign $\{V_{i}(\vec{t})\}_{i=1}^{n}$ into two different groups, which are denoted by $V(R_{1})$ and $V(R_{2})$, where regions $R_{1}$ and $R_{2}$ are visualized in Fig. 1. d, marked by dark blue and light blue respectively. This method is first studied in Ref. [36] to simulate constant 2D digital quantum circuits. From Fig. 1. d, it is shown that each region ($R_{1}$ or $R_{2}$) consists of $\sqrt{n}/4M$ sub-regions which are separated by $\geq 2M$ distance. This property enables operators $V(R_{1})$ and $V(R_{2})$ are easy to simulate classically, and the quantum dynamics mean value has the form $\hat{\mu}(t)=\langle 0^{n}|V(R_{1})V(R_{2})|0^{n}\rangle$. Then the classical Monte Carlo algorithm can be used to approximate $\hat{\mu}(t)$. Noting that operators $V(R_{1})$ and $V(R_{2})$ are not always unitary matrices, they have to be normalized in advance, such that $\gamma_{i}=\|V(R_{i})|0^{n}\rangle\|\leq 1$ for $i\in\{1,2\}$. This step can be implemented efficiently since both $V(R_{1})$ and $V(R_{2})$ are the product of some local operators $V_{i}(\vec{t})$ which can be normalized easily. As a result, as a mean value of $F(x)=\frac{\gamma_{1}\langle x|V(R_{2})|0^{n}\rangle}{\langle x|V^{\dagger}(R_{1})|0^{n}\rangle}$ with $x$ samples from $p(x)=\gamma_{1}^{-1}\left|\langle 0^{n}|V(R_{1})|x\rangle\right|^{2}$, we have
$\hat{\mu}(t)=\sum\limits_{x}\langle 0^{n}|V(R_{1})|x\rangle\langle x|V(R_{2})|0^{n}\rangle=\sum\limits_{x}p(x)\frac{\langle x|V(R_{2})|0^{n}\rangle}{\langle x|V^{\dagger}(R_{1})|0^{n}\rangle},$ (6)
and the variance of $F(x)$ is given by ${\rm Var}(F)=\sum_{x}p(x)\left\|\frac{\gamma_{1}^{2}\langle x|V(R_{2})|0^{n}\rangle}{\langle x|V^{\dagger}(R_{1})|0^{n}\rangle}\right\|^{2}-\hat{\mu}^{2}(t)=\gamma_{1}\gamma_{2}-\hat{\mu}^{2}(t)\leq 1$.
As a result, $\mathcal{O}(4/\epsilon^{2})$ samples $x$ suffice to provide an estimation to $\hat{\mu}(\vec{t})$ within $\mathcal{O}(\epsilon/2)$ additive error. Combining the above two steps together, a $\epsilon$ approximation to the $K$-step quantum mean value problem is provided. We summarize the above steps in Alg. 1. In Appendix D, we provide technical details on how to evaluate $V_{i}(\vec{t})$ and how to compute $F(x)$.
1 Input: Hamiltonian set $\{H^{(1)},\cdots,H^{(K)}\}$, time series $\{t_{1},\cdots,t_{K}\}$, global observable $O$, accuracy $\epsilon$;
2 Output: Mean value estimation $\hat{\mu}(t)$;
3 for $i=1,\cdots,n$
4   Compute $V_{i}(\vec{t})$ given by Eq. 38 via using Lemma 5.
5
6End for
7 Grouping $\{V_{i}(\vec{t})\}$ into $V(R_{1})$ and $V(R_{2})$;
8 for $j=1,\cdots,J=[4/\epsilon^{2}]$
9   Sample $x_{j}\sim p(x)=\left|\langle 0^{n}|V(R_{1})|x\rangle\right|^{2}$, Compute $F(x_{j})$ by using Lemma 7;
10 End for
Output $\hat{\mu}(\vec{t})=\frac{1}{J}\sum_{j=1}^{J}F(x_{j})$.
Algorithm 1 Classical Algorithm for $K$-step Quantum Dynamics Mean Value
V Applications
Let us now turn to applications of our results. We summarize a handful of the most prominent applications below and provide full details in the Appendix.
V.1 Dequantization Quantum Eigenvalue Estimation Algorithm
The guided local Hamiltonian problem plays a significant role in the quantum many-body physics and quantum chemistry simulation. Suppose an initial guided state with $p_{0}\in(1/{\rm poly}(n),1-1/{\rm poly}(n))$ overlap to the quantum ground state, then estimating the ground state energy within an additive error $\delta\leq\mathcal{O}(1/{\rm poly}(n))$ is a BQP-complete problem in the worst-case scenario for certain $2$D Hamiltonians [76]. In this paper, we argue that the guided local $2$D Hamiltonian ground state problem can be classically solved when the accuracy requirement is $\delta\leq\mathcal{O}(1)$. Specifically, given a classical initial state $|\psi_{c}\rangle$ with a $p_{0}\geq\Omega(1/\log n)$ overlap to the target ground state, we theoretically demonstrate the existence of a quasi-polynomial classical algorithm capable of estimating the ground state energy within $\delta\leq\mathcal{O}(1)$ additive error. The fundamental insight arises from symmetry properties enabling the Hadamard test quantum circuit to degenerate to ancilla-free Hadamard test [77] which can be dequantized by using Alg. 1. In particular, we consider the dequantization of the quantum algorithm given by Ref. [48]. Given a $2$D Hamiltonian $H$ with particular symmetry (such as particle number preserving) and a classical initial state $|\psi_{c}\rangle$, sample from the Loschmidt echo $\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle$ can be classically simulated in the running time $\tilde{\mathcal{O}}(n^{e^{t}\log(n)})$. Ref. [48] stated that when the evolution time $t=\mathcal{O}(\log(\delta^{-1}p_{0}^{-1})/\delta)$, a polynomial classical post-processing algorithm may provide a $\delta$-approximation to the ground state energy. Combined all together, the 2D guided local Hamiltonian problem can be classically simulated in quasi-polynomial time when $p_{0}\geq\Omega(1/\log n)$ and constant energy accuracy $\delta$.
Corollary 2.
Given a 2D geometrical local Hamiltonian $H$ that satisfies certain symmetry, and a corresponding classical initial state $|\psi_{c}\rangle$ with $R$ configurations which has $p_{0}$ overlap to the ground state. There exists a classical algorithm that can output $\delta$-approximation to the ground state energy with the run time of
$\mathcal{O}\left(\left(2Rn\right)^{e^{f(p_{0},\delta)}\log(2Rn)+\mathcal{O}(1)}\right),$ (7)
where $f(p_{0},\delta)\leq\mathcal{O}(\delta^{-1}\log(\delta^{-1}p_{0}^{-1}))$.
We further extend this result to the broader and more general problem of Multiple Eigenvalue Estimation (MEE). A typical instance of Multiple Eigenvalue Estimation (MEE) is the estimation of the low-lying energies of Hamiltonian $H$. This has numerous applications, for example, in determining the electronic and optical properties of materials. Similar to the guided local Hamiltonian problem, the MEE problem can also be solved by using samples extracted from the Loschmidt echo $\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle$, where the classical initial state $|\psi_{c}\rangle$ satisfies the “Sufficiently Dominant Condition” assumption [78]. Following the above logic, the quantum MEE algorithm can also be naturally dequantized by using Alg. 1, accompanied by the classical post-processing method given by Ref. [78]. We note that the above approaches are also compatible with the “sampling-access” to the initial state $|\psi_{c}\rangle$, as used in Ref. [37].
Lastly, we emphasize that the above results do not depend on the strong assumption $\|H\|\leq 1$ as used in Ref. [37]. Meanwhile, this section partially answers the open problem mentioned by D. Wild et al. [35] in approximating the Loschmidt echo $\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle$ for $t=\mathcal{O}(1)$. Our result implies that when $H$ and $|\psi_{c}\rangle$ satisfy certain symmetry, the Loschmidt echo can be equivalently transformed into solving a quantum dynamics mean value problem (Problem 1), and the Loschmidt echo thus can be solved by running Alg. 1. In Appendix L, we elaborately detail in simulating the ancilla-free Hadamard test algorithm, meanwhile providing technical details in proving Corollary 2 and solving MEE problems.
V.2 Simulate Superconducting Quantum Computation
In this section, we first demonstrate that classical computers can efficiently simulate algorithms implemented using constant-depth superconducting quantum circuits, applicable to both gate-based and analog superconducting quantum computation. Next, we argue that current noisy superconducting quantum devices may require quasi-polynomial sample complexity to accurately simulate ideal 2D quantum dynamics, especially when factoring in the overhead of quantum error mitigation. Consequently, under these conditions, superconducting quantum computers are unlikely to offer exponential speedup for obtaining expectation values with constant-depth circuits or constant-evolution time.
Definition 1 (Gate-based Superconducting quantum Computation [79]).
Superconducting quantum computers can implement a quantum algorithm by using the elementary gate set $\{e^{-iH_{1q}t},e^{-iH_{2q}t}\}$, with $H_{1q}$ and $H_{2q}$ represent the single- and double-qubit resonator Hamiltonians, respectively, and the evolution time is constrained to $t\leq\mathcal{O}(1)$.
In each quantum layer, the support of involved quantum gates does not have the overlap, as a result, each layer can be represented by a Hamiltonian dynamics $U=e^{-i\sum_{j=1}^{Q}H_{j}t_{j}}$ with $Q\leq\mathcal{O}(n)$, $H_{i}\in\{H_{1q},H_{2q}\}$ and $\left|t_{j}\right|\leq\mathcal{O}(1)$. Given an initial quantum state $|0^{n}\rangle$ and any observable $O=O_{1}\otimes\cdots\otimes O_{n}$, any constant layer superconducting quantum computation can be efficiently simulated by a classical algorithm [36]. We extend this argument to more general scenarios, by endowing much stronger computational capability to superconducting devices.
Definition 2 (Analog Superconducting quantum Computation).
Consider a $\sqrt{n}\times\sqrt{n}$ lattice graph $G=(V,E)$, where vertex set $V$ represents the qubit array and $E$ represents the qubit connection set. The analog superconducting quantum computation can achieve $e^{-iHt}$ in each layer, with $H=\sum_{(i,j)\in E}h_{i,j}$, operator norm $\|h_{i,j}\|\leq 1$ and $t\leq\mathcal{O}(1)$.
As claimed by Theorem 1, we rigorously demonstrate that any $K\leq\mathcal{O}(\log\log n)$-layer analog superconducting quantum computation can be simulated by a classical algorithm with a quasi-polynomial running time in terms of the system size $n$.
On the other hand, due to imperfections in current quantum devices, quantum error mitigation is required to correct the noise-induced bias. As a result, a fair comparison between classical and current quantum computational models should consider the computation cost taken by error mitigation. The basic idea is to correct the effect of quantum noise via classical post-processing on measurement outcomes, without mid-circuit measurements and adaptive gates, as is done in the standard error correction. Here, we argue that the current error mitigation strategies may require a number of samples scaling exponentially in the number of layers. When the quantum circuit depth exceeds $\Omega({\rm poly}\log(n))$, this result thus implies the original quantum advantages may be lost, compared to the proposed classical simulation algorithm (Theorem 1). We extend previous results given by Refs [80, 81] to more general Pauli channels and measurement accuracy, without depending on the unitary $2$-design assumption  [^1] Ref [80] utilized the unitary design property to prove a more compact lower bound, however, it is still unknown whether the “noise channel delay” remains correct for non-Clifford circuits. The error-mitigation overhead can be characterized by the following result.
Theorem 2.
Let $\mathcal{A}$ be an input state-agnostic error mitigation algorithm that takes as input $m$ copies noisy quantum states produced by a $d$-depth quantum circuit affected by $q$-strength local Pauli noise channels, along with a set of observables $\{O\}$. Suppose the algorithm $\mathcal{A}$ is able to produce estimates $\{\hat{o}\}$ such that $\left|\hat{o}-\langle o\rangle\right|\leq\epsilon$. Then the sample complexity
$m\geq\min\left\{\frac{q^{-2cd}(1-\eta)^{2}}{2n},\frac{2^{3n}(1-\epsilon)^{2}}{\epsilon^{2}}\right\}$ (8)
in the worst-case scenario over the choice of the observable set, where $c=1/(2\ln 2)$ and $\eta\in\mathcal{O}(1)$.
Here, we assume each quantum gate is affected by a local Pauli channel which can represent a general noise after engineering. We leave a rigorous definition to Def 12 and proof details to Appendix I.3. When the quantum circuit depth $d\leq\mathcal{O}(\log(n2^{3n}\epsilon^{-2})/\log(1/q))$, the noise strength term dominates the sample complexity, otherwise the sample complexity lower bound grows exponentially with the system size $n$. In the context of superconducting quantum computation, the optimal Hamiltonian simulation algorithm for $2$D local Hamiltonian requires $d\leq\mathcal{O}(t^{2}\log^{3}(nt/\epsilon))$ quantum circuit depth [67], and this thus introduces quasi-polynomial number of samples in the worst-case scenario over the choice of observables.
Furthermore, recent studies have also demonstrated that expectation values on noisy quantum circuits with $\mathcal{O}(\log(n))$ circuit depths are classically simulable [39, 40]. These findings highlight the theoretical limitations of 2D superconducting quantum computation during the NISQ era. However, the proposed classical algorithms are efficient only in the asymptotic sense, with large $n$ and constant error rate $\gamma=1-q$, but still involve large exponents, leaving room for practical quantum advantages. For instance, in intermediate-scale systems with $n=100$ qubits and an error rate $\gamma=10^{-3}\leq 1/n$, the error rate is non-constant, and a circuit depth of 100 is non-constant and generally non-simulable. Moreover, these works assume expectation values without mid-circuit measurements and feedforward operations, which are critical for fault-tolerant quantum computing. Addressing how to classically simulate 2D quantum dynamics without these assumptions remains an intriguing open question.
V.3 Dequantization Quantum Variational Algorithms
Variational quantum algorithms have been experimentally verified for simulating quantum many-body systems and solving combinatorial optimization problems. It is shown that current classical-quantum hybrid workflows can be modeled by a unified approach using a multi-step quantum dynamics mean value problem. For example, consider employing VQE algorithms to approximate the ground state of a 2D Fermi-Hubbard model with dimensions $n_{a}\times n_{b}$. The target ground state is prepared by using a $p$-depth Hamiltonian Variational (HV) ansatz
$\prod_{i=1}^{p}e^{-iH_{v}t_{v}^{(i)}}e^{-iH_{h}t_{h}^{(i)}}e^{-iH_{o}t_{o}^{(i)}},$
where $H_{v},H_{h}$ denote vertical and horizontal hopping terms, respectively, and $H_{o}$ represents the onsite terms [53, 83]. The variational parameters $\{t_{v}^{(i)},t_{h}^{(i)},t_{o}^{(i)}\}_{i=1}^{p}$ are initialized to $[1/p]^{3p}$, and are then refined based on directions determined by classical optimization methods [84]. By applying the superfast encoding method [51, 52, 53], each vertical hopping, horizontal hopping, and onsite term can be encoded into geometrically local Pauli operators with locality $L=5,7,8$, respectively. During the classical optimization process, we have the following result.
Corollary 3.
Given a two-dimensional Fermi-Hubbard model defined on a $(n_{a}\times n_{b})$-sized lattice, a $p$-depth Hamiltonian Variational ansatz with parameters $\{t^{(j)}_{v},t^{(j)}_{h},t^{(j)}_{o}\}_{j=1}^{p}\in[-\pi,\pi]^{3p}$ and a slater determinant initial state, a $\epsilon$-approximation to the VQE energy function can be simulated by a classical algorithm with a run time
$\mathcal{O}\left(\frac{4n_{a}n_{b}}{\epsilon^{2}}\left(\frac{2L}{\epsilon}\right)^{e^{4\pi^{2}ep\mathfrak{d}}\log(2L/\epsilon)}\right),$ (9)
where the constant $\mathfrak{d}$ represents the maximum degree of the interaction graph induced by 2D Fermi-Hubbard model and the locality $L\leq 8$.
It is shown that when $p\sim\mathcal{O}(1)$, the VQE algorithm induced by the 2D Fermi-Hubbard model can be efficiently simulated by a classical algorithm. Furthermore, given the VQE output state $|\psi_{g}\rangle$, Alg. 1 can provide an estimation to $\langle\psi_{g}|O|\psi_{g}\rangle$, where the global observable $O$ may relate to the spin-charge separation, local-gapped phases and other complex topological quantum phases [85, 86, 87, 88] such that the string order parameter [pollmann2012detection] is an exemplary instance for distinguishing symmetry protected topological phases.
Meanwhile, we note that a certain class of QAOA can be classically simulated. Let us consider a MaxCut problem induced by an unweighted $\mathfrak{d}$-regular graph $G_{\rm QAOA}=(V,E)$ with the vertex set $V=(v_{1},\cdots,v_{n})$ and edge set $E=\{e_{ij}\}$. In the context of QAOA, finding the solution can be equivalently achieved by minimizing the loss function $\langle+|^{\otimes n}U^{\dagger}(\vec{\beta},\vec{\gamma})H_{A}U(\vec{\beta},\vec{\gamma})|+\rangle^{\otimes n}$, where the HV ansatz $U(\vec{\beta},\vec{\gamma})=\prod_{k=1}^{p}e^{-i\beta_{k}H_{A}}e^{-i\gamma_{k}H_{B}}$, the problem-oriented Hamiltonian $H_{A}=-\frac{1}{2}\sum_{e_{ij}\in E}(\mathbb{I}-Z_{i}Z_{j})$ and the mixer $H_{B}=\sum_{i=1}^{n}X_{i}$. Since all Pauli terms in $H_{A}$ are local operators, it is interesting to note that such local property enables our algorithm to bypass the $2$D constraint. Specifically, one can directly take $Z_{i}Z_{j}$ into Eq. 38 to substitute local observable $O_{i}$, and $U^{\dagger}(\vec{\beta},\vec{\gamma})Z_{i}Z_{j}U(\vec{\beta},\vec{\gamma})$ can be approximated by an operator $V_{ij}$ such that
$\left\|V_{ij}-U^{\dagger}(\vec{\beta},\vec{\gamma})Z_{i}Z_{j}U(\vec{\beta},\vec{\gamma})\right\|\leq\epsilon,$ (10)
where $V_{ij}$ is essentially a linear combination of ${\rm poly}(n)$ Pauli operators with support size $\leq\mathcal{O}\left(e^{2\pi ep\tau\mathfrak{d}}\log^{2}(1/\epsilon)\right)$, where $\tau=\max\{\left|\beta_{k}\right|,\left|\gamma_{k}\right|\}_{k=1}^{p}$. Ref. [54] numerically benchmarked variational parameter distributions $\vec{\beta},\vec{\gamma}$ for unweighted regular graph MaxCut problem, and they demonstrated that optimal parameters generally living in the interval $[0,0.4\pi]$ which implies $\tau\leq 0.4\pi$. Combining with the classical simulation complexity, we conjecture that a constant-depth QAOA program induced by a constant-regular graph can be efficiently simulated by a classical algorithm. We provide details in Appendixes G, J and K.
V.4 Simulating Adiabatic quantum computation
In this section, we extend to classical simulation of constant-time time-dependent Hamiltonian simulations. The constant-time time-dependent dynamics hold a special interest in many-body physics as they closely relate to the definition of quantum phases. For example, two gapped quantum ground states in the same phase if and only if an adiabatic path with a non-closing (i.e., $\mathcal{O}(1)$) energy gap connects them [62]. This fundamental insight has led to numerous crucial results in studying gapped quantum systems and their phases [89, 63], such as topological orders [64, 90], their stabilities [91] and spectral flow [92, 93] that define quantum phases in the thermodynamics limit.
Consider a system with a one-parameter family of Hamiltonians $H(t):=(1-t)H_{0}+tH_{1}$ for $t\in[0,1]$, where $H_{0}$ is the initial Hamiltonian that adiabatically evolves into the target Hamiltonian $H_{1}$. At any given time $t$, the Hamiltonian $H(t)$ can be expressed as $H(t)=\sum_{X\in S}\lambda_{X}(t)h_{X}$, where $\lambda_{X}(t)$ are time-dependent coefficients and $h_{X}$ are local terms. We assume that a lower-bound estimation $\gamma$ of the energy gap of order $\mathcal{O}(1)$ is known throughout the adiabatic path. Given a local observable $O$ and the adiabatic dynamics described by $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$, where $\mathcal{T}$ denotes the time-ordering operator, we use the estimator $\braket{O(t)}=\bra{\psi}U^{\dagger}(t)OU(t)\ket{\psi}$ for some product state |$\ket{\psi}$. The objective is to provide an estimate $\braket{O}^{\prime}$ that is accurate to within a given $\epsilon$ of the true expectation value. These scenarios are physically relevant as the states that adiabatically connect to a product state are short-range entangled states, which are crucial for studying symmetry-protected topological orders.
To this end, we generalize the conventional cluster expansion to the time-dependent case for simulating short-time adiabatic dynamics. We then apply the cluster expansion for classical simulation of the expectation value of local observables generated from a constant-time adiabatic evolution. Along the path of the derivation of the time-dependent cluster expansion, we also propose a nascent time-dependent Baker–Campbell–Hausdorff (BCH) formula that combines the BCH formula with the Dyson expansion. See Appendix B.4. This may be of independent interest to other quantum science studies, such as the operator dynamics of time-dependent processes. The Heisenberg-picture Dyson series then helps us with the derivation of the time-dependent cluster expansion (See Appendix B.5 for the details). Interestingly, the time-dependent cluster expansion has a similar form to the time-independent one. As such, an algorithm is designed in a similar way to the conventional cluster expansion, which yields the following result (See Appendix M for the proof).
Theorem 3.
Given a family of Hamiltonians $H(t):=(1-t)H_{0}+tH_{1}$ for $t\in[0,1]$ with $\mathfrak{d}$ the maximal degree of the interaction graph, and a local observable $O$, let $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$ be the Hamiltonian evolution operator of the family of Hamiltonians with evolution time $t$. Then, for any $t<t^{*}=\frac{1}{2\sqrt{e\mathfrak{d}}}$, there exists an algorithm with the run time
$\mathrm{poly}\left(\left(\frac{\|O\|}{\epsilon}t^{2}e^{t^{2}}\right)^{\frac{\log\left(\frac{\epsilon}{\|O\|}(t/t^{*}-1)\right)}{\log(t^{*}/t)}}\right)$ (11)
that outputs an estimation $\braket{O}^{\prime}$ to $\braket{O}:=\bra{\psi}U^{\dagger}(t)OU(t)\ket{\psi}$ for some product state |$\ket{\psi}$ within $\epsilon$ accuracy:
$|\braket{O}^{\prime}-\braket{O}|\leq\epsilon.$ (12)
We find that computational complexity experiences a phase-transition-like behavior at $t=t^{*}$ just as the situation encountered in the time-independent case given by Appendix E.1. Comparatively, the transition time is (approximately) quadratically larger for time-dependent dynamics. To lift the simulation time to arbitrary constant time, one may apply analytic continuation. Yet, we find the situation becomes more complicated in the time-dependent scenario as one cannot easily decouple the contribution of real and imaginary parts in the Hamiltonian after applying a mapping function to the time variable due to the non-commutativity of Hamiltonians at different times. We leave this an open question for future exploration.
VI Discussion and Outlook
In this work, we introduce a quasi-polynomial time classical algorithm for simulating constant-time 2D quantum dynamics. We demonstrate its applications in efficiently dequantizing quantum eigenvalues and eigenstates, simulating superconducting quantum computers, dequantizing quantum variational algorithms, and simulating constant-gap adiabatic quantum evolution. Our method and its applications offer innovative approaches to benchmarking short-time 2D fault-tolerant quantum computations and advancing our understanding of 2D quantum system properties, including Hall effects and superconductivity.
This work leaves significant opportunities for further research. Beyond 2D structures, some quantum systems are inherently three-dimensional or higher. A key question is whether our algorithm can be extended to 3D or even higher dimensions. Additionally, concerning the guided local Hamiltonian problem, we have shown that a 2D guided local Hamiltonian problem with a favorable initial state can be solved in linear quasi-polynomial time. Given the quantum PCP conjecture, which suggests that verifying the ground state energy of a local Hamiltonian within constant precision is QMA-hard in the worst case, it remains to be seen whether finding a good initial state is similarly challenging for 2D quantum systems or if the conjecture does not apply in these cases.
Data availability
No datasets were generated or analysed during the current study.
Competing interests
The authors declare no competing interests.
Acknowledgments
The authors would like to express gratitude to Xiaoming Zhang, Jinzhao Sun, Zhongxia Shang, Bujiao Wu and Jingbo B. Wang for valuable discussions. This work is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2023ZD0300200), the National Natural Science Foundation of China Grant (No. 12175003 and No. 12361161602), NSAF (Grant No. U2330201).
Supplementary Information
Appendix A Comparison to related Results
In this section, we review recent progress in studying the classical simulation of short-time Hamiltonian simulation and shallow-depth quantum circuits. It has been long known that the simulation of Hamiltonian evolution or quantum circuits is classically intractable. For example, in the seminal work of Ref. [14], it is shown that the simulation (sample from) of constant-depth quantum circuits is hard unless the polynomial hierarchy collapses. This hardness result persists even for depth-$3$ 2D quantum circuits, indicating the non-simulatability even for constant-depth quantum circuits in low dimensions. Nevertheless, recent efforts have found that generating samples from a constant-depth quantum circuit that supports a geometrical lattice can indeed be simulated efficiently by a classical computer, a counter-intuitive result shows that quantum-classical computational separation may not be fully understood yet. Besides, for more physical-relate purposes, e.g. estimating expectation values of observables with respect to the quantum circuits, the classical simulation could become tractable in certain scenarios [36].
While it is known that Hamiltonian simulation for a polynomial time is a BQP-complete problem, it is still open that whether decreasing the simulation time to constant or poly-logarithmic the problem remains BQP-complete. To this end, Ref. [35] considered the estimation of the expectation value of $k$-local observables $O$ concerning Hamiltonian evolutional circuit: $\langle O(t)\rangle:=\bra{\psi}e^{iHt}Oe^{-iHt}\ket{\psi}$, where |$\ket{\psi}$ is an initial state prescribed to be a product state, $H$ is $\mathfrak{d}$-sparse. Through techniques of cluster expansion [47], which also acts as a key ingredient in this work, a computational complexity scales super-polynomially with $t/t_{c}$ and $\frac{1}{\epsilon}$ is achieved for $t<t_{c}$, where $t_{c}:=1/(2e\mathfrak{d})$ is the critical time for reaching an $\epsilon\|O\|$ error in estimation. Astute readers may find the results surprising as the complexity is independent of the Hamiltonian’s operator norm $\|H\|$, which outperforms the state-of-the-art quantum algorithm [67] that has polynomial-logarithmic dependence on $\|H\|$. The key to this independence is that disconnected clusters have no contribution to the estimator $\langle O(t)\rangle$ resulting from the geometric locality of the Hamiltonian and sparsity of the observable. Furthermore, by devising an ingenious analytic continuation of the estimator $\braket{O(t)}$, Ref. [35] manage to extend the classical simulation to arbitrary $O(1)$ time with the computational cost blow up to scale doubly exponential with $t/t_{c}$, that is $\mathrm{poly}\left((\frac{1}{\epsilon}\frac{t}{t_{c}})^{\frac{t}{t_{c}}}\right)$. However, Ref. [35] can only handle the time evolution of local observable $O$.
On the other hand, the concept of ‘quasi-adiabatic continuation’ (QAC) [94, 95] is developed for extracting physical properties of the adiabatic evolution of constant time. Here, constant-time adiabatic evolution is of special interest because it relates to the definition of quantum phases [62]. The QAC method is built primarily upon the Lieb-Robinson bound [96, 97], which utilizes the fact that in the Heisenberg picture, the short-time dynamics $O(t)=e^{iHt}Oe^{-iHt}$ is (approximately) confined in the light cone supported on the interacting graph of $O(0)$. Yet, when the simulation time remains constant, the Lieb-Robinson bound scales super-polynomially with $e^{\mathcal{O}(vt)}$ and $\frac{1}{\epsilon}$, where $v$ is the Lieb-Robinson velocity. As pointed out in Ref. [35], this result is outperformed by the cluster expansion method with alternatively a polynomial dependence on $\frac{1}{\epsilon}$. While the time-dependence case is not considered in Ref. [35], here we manage to extend to the classical simulation of time-dependent quantum dynamics, where algorithms with similar complexity are achieved.
S. Bravyi et al. [36] proposed a classical algorithm for simulating the quantum mean value problem for general classes of quantum observables $O=O_{1}\otimes\cdots\otimes O_{n}$ and $2$-dimensional constant-depth quantum circuits $U$. Specifically, they divide $UOU^{\dagger}=\prod_{i=1}^{n}UO_{i}U^{\dagger}$ into two operators $U_{A}U_{B}$, where $U_{A}$ and $U_{B}$ can be classically simulated easily. By utilizing the classical Monte Carlo method, they can efficiently simulate the value $\langle 0^{n}|U_{A}U_{B}|0^{n}\rangle$ which approximates the quantum mean value. However, when the unitary $U$ is given by a Hamiltonian dynamics $e^{-iHt}$, the fundamental challenge arises because $e^{-iHt}O_{i}e^{iHt}$ may not be easily computed by the casual principle which only considers quantum gates in the light-cone of $UO_{i}U^{\dagger}$. If we utilize the Trotter-Suzuki method to translate Hamiltonian dynamics $e^{iHt}$ into quantum circuit model $\tilde{U}$, the resulting quantum circuit depth would be ${\rm poly}(\|H\|,t^{O(1)},1/\epsilon^{O(1)})$ [66]. When the Hamiltonian norm $\|H\|=\Theta(n)$ or the accuracy $\epsilon=\mathcal{O}(1/n)$, the corresponding quantum circuit $\tilde{U}$ will spread the information to the whole system, and computing $\tilde{U}O_{i}\tilde{U}^{\dagger}$ by casual principle would be classically hard. In this work, we present a technique able to overcome the above obstacle, by combining the cluster expansion method and analytic continuation. The proposed method can efficiently approximate $e^{-iHt}O_{i}e^{iHt}$ meanwhile limiting its support size to ${\rm poly}\log n$ rather than ${\rm poly}(n)$. As a result, our method nontrivial extends and generalizes Ref. [36] in solving quantum mean value problems when $U$ is given by Hamiltonian dynamics.
From an application perspective, hybrid quantum-classical algorithms, such as quantum approximate optimization algorithms (QAOAs) and variational quantum eigensolvers (VQEs), are paradigmatic protocols for demonstrating the potential quantum advantage on near-term quantum devices. Yet, provable theoretical barriers [98, 99] are found for constant-depth QAOA methods. For solving classical optimization problems, it is discovered that the computational basis measurement of constant-depth QAOA approaches will concentrate regarding the distribution of the output Hamming weight, which can then be used to show their inability to outperform classical algorithms. Our methods on the other hand feature a different direction for classical computation of the output results, which provides a computational-theoretical oriented perspective on the problem.
Appendix B Theoretical Background
B.1 Geometrical Local Hamiltonian
Definition 3 (2D local Hamiltonian).
A two-dimensional Hamiltonian is composed by linear combinations of Hermitian operators $h_{X}$ which nontrivially acts on the qubit subset $X\in S$ with the corresponding coefficient $\lambda_{X}$. Here, the coefficients satisfy $\left|\lambda_{X}\right|\leq 1$ and are chosen such that $\|h_{X}\|=1$. All subsystem set $X\in S$ are geometry local on a two dimensional plane. We define the associated Hamiltonian as $H=\sum_{X\in S}\lambda_{X}h_{X}$.
In this article, we assume Hermitian terms $h_{X}$ are distinct and non-identity multi-qubit Pauli operators. Such assumption naturally satisfies $\|h_{X}\|=1$. computation.
Definition 4 (Operator support [36, 47]).
The support ${\rm supp}(P)$ of an operator $P$ represents the minimal qubit set such that $P=O_{{\rm supp}(P)}\otimes I_{n\setminus{\rm supp}(P)}$ for some operator $O$.
Definition 5 (Superconducting Quantum Computation).
Given a two-dimensional local Hamiltonian $H$, a single-step superconducting quantum computation can be defined by $e^{-iHt}$, where the evolution time $t\leq\mathcal{O}(1)$.
Since the current superconducting quantum chip can only be achieved on a two-dimensional lattice [24], we thus can utilize a two-dimensional local Hamiltonian to characterize the superconducting quantum In the practical superconducting quantum computation, the decoherence time is generally around $10.0$-ns [24] which may not increase with the system size ($n$). As a result, it is reasonable to assume a constant evolution time in our algorithm. Meanwhile, it is interesting to note that the proposed classical algorithm essentially handles a more complex scenario compared to what superconducting quantum computation can really do. In the practical machine, each quantum layer contains several tensor-product quantum gates $g_{1}\otimes\cdots\otimes g_{L}$ which are related to Hamiltonian dynamics $e^{-ih_{1}t_{1}}e^{-ih_{2}t_{2}}\cdots e^{-ih_{L}t_{L}}$. Since $[h_{l},h_{k}]=0$ (each layer can only perform tensor-product gates), the Hamiltonian dynamics can be rewritten by
$e^{-i(h_{1}t_{1}+\cdots h_{L}t_{L})}=e^{-i\left|t_{\max}\right|\sum_{l=1}^{L}\lambda_{l}h_{l}},$ (13)
where $\lambda_{l}=t_{l}/\left|t_{\max}\right|$, $t_{\max}=\max\{t_{l}\}_{l=1}^{L}$ and ${\rm supp}(h_{l})\cap{\operatorname{supp}}(h_{k})=\emptyset$. On the contrary, our classical algorithm can handle a more general case, where each Hermitian term $h_{X}$ has the overlap to at most $\mathfrak{d}\leq\mathcal{O}(1)$ terms. This relationship can be characterized by the cluster and interaction graph model.
B.2 Cluster and Interaction Graph
Definition 6 (Cluster induced by Hamiltonian).
Given a general two-dimensional Hamiltonian
$H=\sum_{X\in S}\lambda_{X}h_{X},$
a cluster $\bm{V}$ is defined as a nonempty multi-set of subsystems from $S$, where multi-sets allow an element appearing multiple times. The set of all clusters $\bm{V}$ with size $m$ is denoted by $\mathcal{C}_{m}$ and the set of all clusters is represented by $\mathcal{C}=\cup_{m\geq 1}\mathcal{C}_{m}$.
For example, if the Hamiltonian $H=X_{0}X_{1}+Y_{0}Y_{1}$, then some possible candidates for $\bm{V}$ would be $\{X_{0}X_{1}\}$, $\{Y_{0}Y_{1}\}$, $\{X_{0}X_{1},X_{0}X_{1}\},\cdots$. We call the number of times a subsystem $X$ appears in a cluster $\bm{V}$ the multiplicity $\mu_{\bm{V}}(X)$, otherwise we assign $\mu_{\bm{V}}(X)=0$. Traversing all subsets $X\in S$ may determine the size of $\bm{V}$, that is $\left|\bm{V}\right|=\sum_{X\in S}\mu_{\bm{V}}(X)$. In the provided example, when $\bm{V}=\{X_{0}X_{1},X_{0}X_{1}\}$, we have $\mu_{\bm{V}}(X_{0}X_{1})=2$, $\mu_{\bm{V}}(Y_{0}Y_{1})=0$ and $\left|\bm{V}\right|=2$.
Definition 7 (Interaction Graph).
We associate with every cluster $\bm{V}$ a simple graph $G_{\bm{V}}$ which is also termed as the cluster graph. The vertices of $G_{\bm{V}}$ correspond to the subsystems in $\bm{V}$, with repeated subsystems also appearing as repeated vertices. Two distinct vertices $X$ and $Y$ are connected by an edge if and only if the respective subsystems overlap, that is ${\rm supp}(h_{X})\cap{\rm supp}(h_{Y})\neq\emptyset$.
Suppose the cluster $\bm{V}=\{X_{0}X_{1},X_{0}X_{1}\}$, then its corresponding interaction graph $G_{\bm{V}}$ has two vertices $v_{1},v_{2}$, related to $X_{0}X_{1}$ and $X_{0}X_{1}$, respectively, and $v_{1}$ connects to $v_{2}$ since ${\rm supp}(X_{0}X_{1})\cap{\rm supp}(X_{0}X_{1})\neq\emptyset$. We say a cluster $\bm{V}$ is connected if and only if $G_{\bm{V}}$ is connected. We use the notation $\mathcal{G}_{m}$ to represent all connected clusters of size $m$ and $\mathcal{G}=\cup_{m\geq 1}\mathcal{G}_{m}$ for the set of all connected clusters.
Definition 8 (Super-Interaction Graph).
Suppose we have $K$ clusters $\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K}$, we define the super-interaction graph $G^{K}_{\bm{V}_{1},\cdots,\bm{V}_{K}}$ composed by interaction graphs $G_{\bm{V}_{1}},G_{\bm{V}_{2}},\cdots G_{\bm{V}_{K}}$, where vertices $\{h_{X}\}_{X\in S_{1}\cup S_{2}\cdots\cup S_{K}}$ inherit from $G_{\bm{V}_{1}},G_{\bm{V}_{2}},\cdots G_{\bm{V}_{K}}$ and vertices $h_{X}$ and $h_{Y}$ are connected if ${\rm supp}(h_{X})\cap{\rm supp}(h_{Y})\neq\emptyset$.
In our paper, the super-interaction graph is generally induced by a Hamiltonian series, say $\{H^{(1)},\cdots,H^{(K)}\}$. From the above definition, we know that the super-interaction graph $G^{K}_{\bm{V}_{1},\cdots,\bm{V}_{K}}$ contains $\sum_{k=1}^{K}\left|G_{\bm{V}_{k}}\right|$ vertices.
Definition 9 (Connected Super-Interaction Graph).
The super-interaction-graph $G^{K}_{\bm{V}_{1},\cdots,\bm{V}_{K}}$ is connected if and only if the super cluster $\bm{V}=(\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K})$ is connected. All $m$-sized connected super-interaction graphs are denoted by $\mathcal{G}_{m}^{K}$, with $m=\sum_{k=1}^{K}\left|\bm{V}_{k}\right|$.
Specifically, we denote $\mathcal{G}_{m}^{K,O_{i}}$ as the set of all $m$-sized connected super-interaction graphs which connects to $O_{i}$.
B.3 Cluster Expansion
We first consider a simple case, that is the cluster expansion of the single-step Hamiltonian dynamics $e^{iHt}O_{i}e^{-iHt}$ [35]. For any cluster $\bm{V}\in\mathcal{C}_{m}$, we can write $\bm{V}=(X_{1},\cdots,X_{m})$. This notation helps us to write the function $e^{iHt}O_{i}e^{-iHt}$ as the multivariate Taylor-series expansion by using the cluster expansion method. Here, we fix the parameter $O_{i}$, but considering $\{t,\lambda_{X}\}$ as variables. As a result, we have
$\begin{split}e^{iHt}O_{i}e^{-iHt}=&\sum\limits_{m=0}^{+\infty}\frac{t^{m}}{m!}\left(\frac{\partial^{m}[e^{iHt}O_{i}e^{-iHt}]}{\partial t^{m}}\right)_{t=0}\end{split}$ (14)
Recall that the Hamiltonian $H=\sum_{X\in S}\lambda_{X}h_{X}$, then we assign $z_{X}=-it\lambda_{X}$. This results in
$\frac{\partial[e^{iHt}O_{i}e^{-iHt}]}{\partial t}=\sum_{X\in S}\frac{\partial[e^{iHt}O_{i}e^{-iHt}]}{\partial z_{X}}\frac{\partial z_{X}}{\partial t}=\sum_{X\in S}(-i)\lambda_{X}\frac{\partial[e^{iHt}O_{i}e^{-iHt}]}{\partial z_{X}}.$ (15)
Taking above derivative function into Eq. 14, we have
$\begin{split}e^{iHt}O_{i}e^{-iHt}=&\sum\limits_{m=0}^{+\infty}\frac{(-it)^{m}}{m!}\sum\limits_{X_{1},\cdots,X_{m}}\lambda_{X_{1}}\cdots\lambda_{X_{m}}\left(\frac{\partial^{m}[e^{iHt}O_{i}e^{-iHt}]}{\partial z_{X_{1}}\cdots\partial z_{X_{m}}}\right)_{z=(0,\cdots,0)}\\ =&\sum\limits_{m=0}^{+\infty}(-it)^{m}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\bm{\lambda}^{\bm{V}}}{\bm{V}!}\left(\frac{\partial^{m}[e^{iHt}O_{i}e^{-iHt}]}{\partial z_{X_{1}}\cdots\partial z_{X_{m}}}\right)_{z=(0,\cdots,0)}\end{split}$ (16)
where $\bm{\lambda}^{\bm{V}}=\prod_{X\in S}\lambda_{X}^{\mu_{\bm{V}}(X)}$ and $\bm{V}!=\prod_{X\in S}\mu_{\bm{V}}(X)!$. Finally, we utilize the BCH expansion to compute
$\begin{split}\left(\frac{\partial^{m}[e^{iHt}O_{i}e^{-iHt}]}{\partial z_{X_{1}}\cdots\partial z_{X_{m}}}\right)_{z=(0,\cdots,0)}=&\frac{\partial^{m}}{\partial z_{X_{1}}\cdots\partial z_{X_{m}}}\sum\limits_{j=0}^{\infty}\frac{(-it)^{j}}{j!}[H,O_{i}]_{j}\big{|}_{z=(0,\cdots,0)}\\ =&\frac{(-it)^{m}}{m!}\frac{\partial^{m}}{\partial z_{X_{1}}\cdots\partial z_{X_{m}}}[\underbrace{H,[H,\cdots[H}_{m},O_{i}]\cdots]]\big{|}_{z=(0,\cdots,0)}\\ =&\frac{(-it)^{m}}{m!}\sum_{\sigma\in\mathcal{P}_{m}}[\partial_{z_{X_{\sigma(1)}}}H,\cdots[\partial_{z_{X_{\sigma(m)}}}H,O_{i}]\cdots]\big{|}_{z=(0,\cdots,0)}\\ =&\frac{1}{m!}\sum_{\sigma\in\mathcal{P}_{m}}[h_{X_{\sigma(1)}},\cdots[h_{X_{\sigma(m)}},O_{i}]\cdots].\end{split}$ (17)
As a result, the cluster expansion of the single-step Hamiltonian dynamics can be written by
$e^{iHt}O_{i}e^{-iHt}=\sum_{m\geq 0}^{+\infty}\sum_{\bm{V}\in\mathcal{C}_{m}}\frac{\bm{\lambda}^{\bm{V}}}{\bm{V}!}\frac{(-it)^{m}}{m!}\sum\limits_{\sigma\in\mathcal{P}_{m}}\left[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m)}},O_{i}]\right].$ (18)
Here $\mathcal{P}_{m}$ represents the permutation group on the set $\{1,\cdots,m\}$. We denote the cluster derivative
$D_{\bm{V}}[e^{iHt}O_{i}e^{-iHt}]=\frac{(-it)^{m}}{m!}\sum\limits_{\sigma\in\mathcal{P}_{m}}\left[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m)}},O_{i}]\right].$ (19)
From Eq. 19, we know that $V_{\sigma(1)}\cap V_{\sigma(2)}\cap\cdots\cap V_{\sigma(m)}\cap{\rm supp}(O_{i})=\emptyset$ may result in $D_{\bm{V}}\left[e^{iHt}O_{i}e^{-iHt}\right]=0$ [47, 35]. This property dramatically reduces the computational complexity in approximating $e^{iHt}O_{i}e^{-iHt}$, which only needs to consider connected clusters $\bm{V}$ with bounded size.
B.4 Heisenberg-picture Dyson series
In this section, we provide the derivation of the Heisenberg-picture Dyson series. Given a time-dependent Hamiltonian evolution operator $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$ and an operator $O$, we are interested in expanding the time-evolved operator $O(t)=U^{\dagger}(t)OU(t)$ in the Heisenberg picture using the BCH formula. To tackle the time dependency, a standard technique is the Dyson series. We begin by reviewing the properties of the Dyson series, which gives
$\mathcal{T}\left[e^{-i\int_{0}^{t}H(s)\mathrm{d}s}\right]=\sum_{k=0}^{\infty}\frac{(-i)^{k}}{k!}\int_{0}^{t}\cdots\int_{0}^{t}\mathcal{T}\left[H\left(t_{k}\right)\cdots H\left(t_{1}\right)\right]\mathrm{d}^{k}t,$ (20)
where $\mathcal{T}$ is the time-ordering operator. The time-ordering operator will put time-dependent operators in non-decreasing order according to the time variables, i.e., $\mathcal{T}(H(t_{k})\cdots H(t_{2})H(t_{1}))=H(t_{\sigma(k)})\cdots H(t_{\sigma(2)})H(t_{\sigma(1)})$ such that $\sigma$ is a permutation satisfying $t_{\sigma(1)}\leq t_{\sigma(2)}\leq\cdots\leq t_{\sigma(k)}$. Following Ref. [100], we can discretize the integrals for Riemann integrable $H(t)$ as
$\mathcal{T}\left[e^{-i\int_{0}^{t}H(s)\mathrm{d}s}\right]=\lim_{M\rightarrow\infty}\sum_{k=0}^{\infty}\frac{(-it)^{k}}{k!M^{k}}\tilde{B}_{k},\quad\tilde{B}_{k}=\sum_{m_{1},\cdots,m_{k}=0}^{M-1}\mathcal{T}\left[H\left(m_{k}\Delta\right)\cdots H\left(m_{1}\Delta\right)\right],$ (21)
where $\Delta=t/M$.
At first glance, as we need to resort to the Dyson series for approximating $U(t)$, it seems challenging to expand $O(t)$ in the Heisenberg picture using the BCH formula for the derivation of the cluster expansion as what is done in the time-independent case (see Appendix B.3). Here, we propose the first Heisenberg-picture Dyson series, i.e., combined with the discretized BCH formula. To this end, we mildly generalize the definition of the time-ordering operator $\mathcal{T}$. That is when acting on a product of operators both time-dependent and time-dependent ones, $\mathcal{T}$ will independently act on each consecutive segment of time-dependent operators and leave the time-independent one out. For instance, $\mathcal{T}(H(t_{a_{k}})\cdots H(t_{a_{2}})H(t_{a_{2}})PH(t_{b_{l}})\cdots H(t_{b_{2}})H(t_{b_{2}}))=\mathcal{T}(H(t_{a_{k}})\cdots H(t_{a_{2}})H(t_{a_{2}}))P\mathcal{T}(H(t_{b_{l}})\cdots H(t_{b_{2}})H(t_{b_{2}}))$, where $P$ is some time-independent operator.
Using the generalized time-ordering operator, we define
$\tilde{O}_{r,s}:=\sum_{m_{1},\cdots,m_{r};z_{1},\cdots,z_{s}=0}^{M-1}\mathcal{T}[H_{m_{r}}\cdots H_{m_{1}}OH_{z_{s}}\cdots H_{z_{1}}]=\sum_{m_{1},\cdots,m_{r};z_{1},\cdots,z_{s}=0}^{M-1}\mathcal{T}[H_{m_{r}}\cdots H_{m_{1}}]O\mathcal{T}[H_{z_{s}}\cdots H_{z_{1}}],$ (22)
where we have used the abbreviation $H(m_{r}\Delta)$ for $H_{m_{r}}$, and the second equation is due to our generalization of the time-ordering operator. Besides, we use the adjoint notation to denote the commutator as $\mathrm{ad}_{X}(Y)=[X,Y]$. We can then write the nested commutator of time-dependent Hamiltonians into the form:
$\mathcal{T}[\mathrm{ad}^{j}_{H}(O)]:=\mathcal{T}\left[\sum_{m_{j},\cdots,m_{1}=0}^{M-1}\left[H_{m_{j}},\cdots,\left[H_{m_{1}},O\right]\cdots\right]\right].$ (23)
We propose a helpful lemma for the nested commutator to deduce the BCH formula.
Lemma 1.
Let $\mathcal{T}[\mathrm{ad}^{j}_{H}(O)]$ be a nested commutator defined by Eq. (23). Then, we have
$\mathcal{T}[\mathrm{ad}^{j}_{H}(O)]=\sum_{i=0}^{j}(-1)^{i}\binom{j}{i}\tilde{O}_{j-i,i},$ (24)
where $\tilde{O}_{j-i,i}$ is given by Eq. (22).
Proof.
This is proved by induction. First, we inspect that the first-order expansion is valid for our formula. Then, we proceed by assuming that $j$-th expansion is feasible and check the $(j+1)$-th term:
$\mathcal{T}[\mathrm{ad}^{j+1}_{H}(O)]$ $=\sum_{i=0}^{j}(-1)^{i}\binom{j}{i}\widetilde{[H_{j+1},O]}_{j-i,i}$ (25)
$=\sum_{i=0}^{j}(-1)^{i}\binom{j}{i}\tilde{O}_{j+1-i,i}-\sum_{i=0}^{j}(-1)^{i}\binom{j}{i}\tilde{O}_{j-i,i+1}$
$=\tilde{O}_{j+1,0}+\sum_{i=1}^{j}(-1)^{i}\binom{j}{i}\tilde{O}_{j+1-i,i}-\sum_{i^{\prime}=1}^{j}(-1)^{i^{\prime}-1}\binom{j}{i^{\prime}-1}\tilde{O}_{j+1-i^{\prime},i^{\prime}}+(-1)^{j+1}\tilde{O}_{0,j+1}$
$=\sum_{i=0}^{j+1}(-1)^{i}\binom{j+1}{i}\tilde{O}_{j+1-i,i},$
where in the third line we have substituted $i$ into $i^{\prime}=i+1$ for the second summation in the second line; and in the last line we have used that fact that $\binom{j}{i}+\binom{j}{i-1}=\binom{j+1}{i}$. ∎
We then apply the discretized Dyson expansion provided in Eq. (21) to the Heisenberg picture formula, and thereby obtain the following Heisenberg-picture (discretized) Dyson series, which is summarized as
Lemma 2.
Let $O(t)=\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$ be an observable evolved by a time-dependent evolution in the Heisenberg picture. Then, we can expand $O(t)$ as
$O(t)=\lim_{M\rightarrow\infty}\sum_{j=0}^{\infty}\frac{(it)^{j}}{j!M^{j}}\mathcal{T}[\mathrm{ad}_{H}^{j}(O)].$ (26)
Proof.
We prove by simply applying the result given by Lemma 1:
$O(t)$ $=\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$ (27)
$=\lim_{M\rightarrow\infty}\sum_{p,q=0}^{\infty}(-1)^{q}\frac{(it)^{p+q}}{p!q!M^{p+q}}\tilde{B}_{p}O\tilde{B}_{q}$
$=\lim_{M\rightarrow\infty}\sum_{s}^{\infty}\sum_{d=0}^{s}(-1)^{d}\left(\frac{it}{M}\right)^{s}\frac{\tilde{O}_{s-d,d}}{(s-d)!d!}=\lim_{M\rightarrow\infty}\sum_{j=0}^{\infty}\frac{(it)^{j}}{j!M^{j}}\mathcal{T}[\mathrm{ad}_{H}^{j}(O)].$
Here, in the second line, we insert the discretized Dyson expansion. We then substitute $p,q$ into $s-d,d$ and also note that $\tilde{O}_{s-d,d}=\tilde{B}_{s-d}O\tilde{B}_{d}$ as prescribed by Eq. (22). Eventually, we obtain the final result by applying the conclusion on the nested commutator given by Lemma 1. ∎
B.5 Cluster expansion of time-dependent Hamiltonian dynamics
In this section, we deduce cluster expansion of time-dependent Hamiltonian dynamics following the ideas from Sec. B.3 with a special interest in adiabatic evolution. Our aim is to derive the multi-variate Taylor expansion for $O(t)=U^{\dagger}(t)OU(t)$ with $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$. We also assume that the derivative of $H(t)$ higher than order $1$ vanishes, which is reasonable for typical adiabatic dynamics. We note that for each term in the time-dependent Hamiltonian, the time-dependent part resides in the coefficients.
Theorem 4 (Time-dependent cluster expansion).
Given $O(t)=\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}$, its cluster expansion is
$\begin{split}O(t)=\lim_{M\rightarrow\infty}\sum\limits_{m=0}^{+\infty}\frac{(-it)^{m}}{m!M^{m}}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\tilde{\bm{\lambda}}^{\bm{V}}}{\bm{V}!}&\Bigg{(}\mathcal{T}\Bigg{[}\sum_{n_{m},\cdots,n_{1}=0}^{M-1}\sum_{\sigma\in\mathcal{P}_{m}}\Big{[}h_{X_{\sigma(m)}}f(n_{m},t,X_{\sigma(m)}),\cdots,\\ &\left[h_{X_{\sigma(1)}}f(n_{1},t,X_{\sigma(1)}),O\Big{]},\cdots\right]\Bigg{]}\Bigg{)}_{z=(0,\cdots,0)},\end{split}$ (28)
where $\tilde{\lambda}_{X}(t)=\lambda_{X}(t)+t\lambda^{\prime}_{X}(t)$ with $\lambda^{\prime}_{X}(t)=\frac{\mathrm{d}\lambda_{X}(t)}{\mathrm{d}t}$, $\tilde{\bm{\lambda}}^{\bm{V}}=\prod_{X\in S}\tilde{\lambda}_{X}^{\mu_{\bm{V}}(X)}$, $\bm{V}!=\prod_{X\in S}\mu_{\bm{V}}(X)!$, $f(n_{m},t,X):=\frac{\partial Z_{X}(n_{m})}{\partial Z_{X}(t)}$ and $Z_{X}(t)=-it\lambda_{X}(t)$.
Especially, as we consider the adiabatic Hamiltonian of the form $H(t)=(1-t)H_{0}+tH_{1}$, we could write the Hamiltonian as a sum of local operators $H(t)=\sum_{X\in S}\lambda_{X}(t)h_{X}$. We have the following as the Taylor expansion of $O(t)$ with $\{t,\lambda_{X}(t)\}$:
$\begin{split}O(t)=&\sum\limits_{m=0}^{+\infty}\frac{t^{m}}{m!}\left(\frac{\partial^{m}[\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}]}{\partial t^{m}}\right)_{t=0}.\end{split}$ (29)
Then, by changing of variable as $Z_{X}(t)=-it\lambda_{X}(t)$, we have
$\frac{\mathrm{d}}{\mathrm{d}t}O(t)=\sum_{X}-i\tilde{\lambda}_{X}(t)\frac{\partial[\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}]}{\partial Z_{X}(t)},$ (30)
where $\tilde{\lambda}_{X}(t)=\lambda_{X}(t)+t\lambda^{\prime}_{X}(t)$ with $\lambda^{\prime}_{X}(t)=\frac{\mathrm{d}\lambda_{X}(t)}{\mathrm{d}t}$. Similarly, we take the above result to Eq. (29), which results in
$\begin{split}O(t)=&\sum\limits_{m=0}^{+\infty}\frac{(-it)^{m}}{m!}\sum\limits_{X_{1},\cdots,X_{m}}\tilde{\lambda}_{X_{1}}(t)\cdots\tilde{\lambda}_{X_{m}}(t)\left(\frac{\partial^{m}[\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}]}{\partial z_{X_{1}}(t)\cdots\partial z_{X_{m}}(t)}\right)_{z=(0,\cdots,0)}\\ =&\sum\limits_{m=0}^{+\infty}(-it)^{m}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\tilde{\bm{\lambda}}^{\bm{V}}}{\bm{V}!}\left(\frac{\partial^{m}[\mathcal{T}e^{i\int_{0}^{t}H(s)\mathrm{d}s}O\mathcal{T}e^{-i\int_{0}^{t}H(s)\mathrm{d}s}]}{\partial z_{X_{1}}(t)\cdots\partial z_{X_{m}}(t)}\right)_{z=(0,\cdots,0)}\\ =&\lim_{M\rightarrow\infty}\sum\limits_{m=0}^{+\infty}(-it)^{m}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\tilde{\bm{\lambda}}^{\bm{V}}}{\bm{V}!}\left(\frac{\partial^{m}}{\partial z_{X_{1}}(t)\cdots\partial z_{X_{m}}(t)}\sum_{j=0}^{\infty}\frac{(it)^{j}}{j!M^{j}}\mathcal{T}[\mathrm{ad}_{H}^{j}(O)]\right)_{z=(0,\cdots,0)}\\ =&\lim_{M\rightarrow\infty}\sum\limits_{m=0}^{+\infty}\frac{t^{2m}}{m!M^{m}}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\tilde{\bm{\lambda}}^{\bm{V}}}{\bm{V}!}\left(\mathcal{T}\left[\sum_{n_{m},\cdots,n_{1}}\sum_{\sigma\in\mathcal{P}_{m}}\left[\frac{\partial H_{n_{m}}}{\partial Z_{X_{\sigma(m)}}(t)},\cdots,\left[\frac{\partial H_{n_{1}}}{\partial Z_{X_{\sigma(m)}}(t)},O\right],\cdots\right]\right]\right)_{z=(0,\cdots,0)}.\end{split}$ (31)
Here, in the second line, we apply $\tilde{\bm{\lambda}}^{\bm{V}}=\prod_{X\in S}\tilde{\lambda}_{X}^{\mu_{\bm{V}}(X)}$ and $\bm{V}!=\prod_{X\in S}\mu_{\bm{V}}(X)!$. Then, we substitute in the BCH expansion of $O(t)$ that is derived by Eq. (26). In this last line, we note that the only non-zero term in the summation over $j$ is $j=m$. This is because, for $j<m$ ones, the derivative vanishes because of our assumption about the derivative of the Hamiltonian; for $j>m$ ones, there will be some Hamiltonian left untouched by the partial differentiation operation, which also vanishes because when $Z=0$, the corresponding Hamiltonian becomes zero.
Then, we remark that because each $H_{n_{m}}=(1-\frac{n_{m}t}{M})H_{0}+\frac{n_{m}t}{M}H_{1}=\sum_{X}\lambda_{X}(n_{m})h_{X}$, we have
$\frac{\partial H_{n_{m}}}{\partial Z_{X_{\sigma(m)}}(t)}=\frac{\partial H_{n_{m}}}{\partial Z_{X_{\sigma(m)}}(n_{m})}\cdot\frac{\partial Z_{X_{\sigma(m)}}(n_{m})}{\partial Z_{X_{\sigma(m)}}(t)}=\frac{h_{X_{\sigma(m)}}}{it}f(n_{m},t,X_{\sigma(m)}),$ (32)
where we have taken the abbreviation $f(n_{m},t,X):=\frac{\partial Z_{X}(n_{m})}{\partial Z_{X}(t)}$. Taking this result to Eq. (31), we finally obtain
$\begin{split}O(t)=\lim_{M\rightarrow\infty}\sum\limits_{m=0}^{+\infty}\frac{(-it)^{m}}{m!M^{m}}\sum\limits_{\bm{V}\in\mathcal{C}_{m},V=(X_{1},\cdots,X_{m})}\frac{\tilde{\bm{\lambda}}^{\bm{V}}}{\bm{V}!}&\Bigg{(}\mathcal{T}\Bigg{[}\sum_{n_{m},\cdots,n_{1}}\sum_{\sigma\in\mathcal{P}_{m}}\Big{[}h_{X_{\sigma(m)}}f(n_{m},t,X_{\sigma(m)}),\cdots,\\ &\left[h_{X_{\sigma(1)}}f(n_{1},t,X_{\sigma(1)}),O\Big{]},\cdots\right]\Bigg{]}\Bigg{)}_{z=(0,\cdots,0)}.\end{split}$ (33)
Appendix C Cluster expansion of $K$-step Hamiltonian dynamics
In this article, we consider the $K$-step scenario driven by $\{H^{(1)},\cdots,H^{(K)}\}$ and corresponding time parameters $\{t_{1},\cdots,t_{K}\}$. According to the linear property of the commute net, for any Hermitian operator $A$, we have
$\left[A,\sum\limits_{\sigma\in\mathcal{P}_{m}}[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m)}},O_{i}]]\right]=\sum\limits_{\sigma\in\mathcal{P}_{m}}\left[A,[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m)}},O_{i}]]\right].$ (34)
We first consider the cluster expansion of $2$-step Hamiltonian dynamics
$\begin{split}&e^{iH^{(2)}t_{2}}e^{iH^{(1)}t_{1}}O_{i}e^{-iH^{(1)}t_{1}}e^{-iH^{(2)}t_{2}}\\ =&\sum\limits_{m_{2}\geq 0}\sum\limits_{\bm{V}_{2}\in\mathcal{C}_{m_{2}}}\frac{\bm{\lambda}^{\bm{V}_{2}}}{\bm{V}_{2}!}D_{\bm{V}_{2}}\left[e^{iH^{(2)}t_{2}}\sum_{m_{1}\geq 0}^{+\infty}\sum_{\bm{V}_{1}\in\mathcal{C}_{m_{1}}}\frac{\bm{\lambda}_{1}^{\bm{V}_{1}}}{\bm{V}_{1}!}\frac{(-it)^{m_{1}}}{m_{1}!}\sum\limits_{\sigma\in\mathcal{P}_{m_{1}}}\left[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m_{1})}},O_{i}]\right]e^{-iH^{(2)}t_{2}}\right]\\ =&\sum\limits_{m_{2}\geq 0}\sum\limits_{\bm{V}_{2}\in\mathcal{C}_{m_{2}}}\frac{\bm{\lambda}^{\bm{V}_{2}}}{\bm{V}_{2}!}\frac{(-it_{2})^{m_{2}}}{m_{2}!}\sum\limits_{\sigma_{2}\in\mathcal{P}_{m_{2}}}\left[h_{V_{\sigma_{2}(1)}}\cdots\left[h_{V_{\sigma_{2}(m_{2})}},\sum_{m_{1}\geq 0}^{+\infty}\sum_{\bm{V}_{1}\in\mathcal{C}_{m_{1}}}\frac{\bm{\lambda}_{1}^{\bm{V}_{1}}}{\bm{V}_{1}!}\frac{(-it)^{m_{1}}}{m_{1}!}\sum\limits_{\sigma\in\mathcal{P}_{m_{1}}}\left[h_{V_{\sigma(1)}},\cdots[h_{V_{\sigma(m_{1})}},O_{i}]\right]\right]\right]\\ =&\sum\limits_{m_{1},m_{2}\geq 0}\sum\limits_{(\bm{V}_{1},\bm{V}_{2})}\frac{\bm{\lambda}^{\bm{V}_{1}}\bm{\lambda}^{\bm{V}_{2}}}{\bm{V}_{1}!\bm{V}_{2}!}\frac{(-it_{1})^{m_{1}}(-it_{2})^{m_{2}}}{m_{1}m_{2}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \sigma_{2}\in\mathcal{P}_{m_{2}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots\left[h_{V_{\sigma_{1}(m_{1})}}\cdots\left[h_{V_{\sigma_{K}(m_{2})}},O_{i}\right]\right]\right],\end{split}$ (35)
where the second equality comes from the relationship given by Eq. 34. Repeat above process for $K$ times, we have the cluster expansion of $K$-step Hamiltonian dynamics, that is
$U_{i}(\vec{t})=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}\sum\limits_{\begin{subarray}{c}\bm{V}_{1}\in\mathcal{C}_{m_{1}}\\ \cdots\\ \bm{V}_{K}\in\mathcal{C}_{m_{K}}\end{subarray}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots\left[h_{V_{\sigma_{1}(m_{1})}},\cdots\left[h_{V_{\sigma_{K}(m_{K})}},O_{i}\right]\right]\right].$ (36)
Here, notations $\sigma_{1},\cdots,\sigma_{K}$ represent $K$ permutations, and $\mathcal{P}_{m_{1}}\cdots,\mathcal{P}_{m_{K}}$ represents corresponding permutation groups.
Similar to the single-step Hamiltonian dynamics, we know that if clusters $\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K}$ and $O_{i}$ are disconnected, then the commute net $\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]\right]=0$, which can be summrized as the following lemma.
Lemma 3.
Given clusters $\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K}$ and an observable $O_{i}$, if the supper-interaction graph induced by $\bm{V}=(\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K},O_{i})$ is disconnected, then the commute net
$\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{1}(m_{1})}}\cdots[h_{V_{\sigma_{K}(1)}}\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]]]\right]=0,$
where $\left|\bm{V}_{k}\right|=m_{k}$ and $\sigma_{k}(1),\sigma_{k}(2),\cdots,\sigma_{k}(m_{k})$ represents an entry of the permutation group $\mathcal{P}_{m_{k}}$.
Proof.
Denote all connected super-interaction graph as $\mathcal{G}_{\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K},O_{i}}^{K}$. Consider a cluster $\bm{W}\notin\mathcal{G}_{\bm{V}_{1},\bm{V}_{2},\cdots,\bm{V}_{K},O_{i}}^{K}$. For every permutation series $(\sigma_{1}(1),\cdots,\sigma_{1}(m_{1}),\cdots,\sigma_{K}(m_{K}))$, there exists an index $\sigma_{k}(s)$ such that $\bm{W}_{\sigma_{k}(s)}$ and $\bm{W}_{\sigma_{k}(s+1)}\cup\cdots\cup\bm{W}_{\sigma_{K}(m_{K})}\cup{\rm supp}(O_{i})$ does not have an overlap. This directly results in
$\left[h_{W_{\sigma_{k(s)}}},\cdots[h_{W_{\sigma_{k(m_{k})}}}\cdots[h_{W_{\sigma_{K}(1)}}\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]]]\right]=0,$ (37)
and the concerned commutator vanishes. ∎
Using this property, we may rewrite the above expression by introducing the connected cluster set $\mathcal{G}_{m}^{K,O_{i}}$ composed by all connected super-interaction graphs $G^{K}_{\bm{V}_{1},\cdots,\bm{V}_{K}}$ (connected to $O_{i}$) with size
$m=\left|\bm{V}_{1}\right|+\cdots+\left|\bm{V}_{K}\right|.$
Here, $O_{i}$ is a single-qubit operator non-trivially performs on qubit $i$, then $\{\bm{V}_{1},\cdots,\bm{V}_{K},O_{i}\}$ are connected implies ${\rm supp}(O_{i})\in\bm{V}_{k}$ for some $k\in[K]$. Such observation enables us to only consider summation over $\mathcal{G}_{m}^{K,O_{i}}$, meanwhile truncate the cluster expansion up to $M$ order, that is
$V_{i}(\vec{t})=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}^{M}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]\right].$ (38)
Given above knowledge, we can outline our algorithm and provide the running time complexity analysis.
Appendix D Methods outline and Algorithm Complexity Analysis
D.1 Approximate $V_{i}(\vec{t})$
The following lemma proves the convergence of the $K$-step cluster expansion and the support of $V_{i}(\vec{t})$ for constant times.
Lemma 4 (Informal).
Given a single qubit observable $O_{i}$, then for any $K$-step quantum dynamics driven by $\{H^{(1)},\cdots,H^{(K)}\}$ and corresponding constant time parameters $\{t_{1},\cdots,t_{K}\}$, the operator $U_{i}(\vec{t})=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}}$ can be approximated by an operator $V_{i}(\vec{t})$ such that $\|U_{i}(\vec{t})-V_{i}(\vec{t})\|\leq\epsilon/2L$. Here, $V_{i}(\vec{t})$ represents a $M=\tilde{\mathcal{O}}\left(e^{\pi teK\mathfrak{d}}\log(2L/\epsilon)\right)$-order truncated cluster expansion given by Eq. 38, $\mathfrak{d}$ represents the maximum degree of interaction graphs induced by Hamiltonians $\{H^{(1)},\cdots,H^{(K)}\}$ and $t=\max\{t_{k}\}$. Meanwhile, we have
$\Omega(M)\leq{\rm supp}(V_{i}(\vec{t}))\leq\mathcal{O}(4M^{2}).$ (39)
We leave proof details in Appendix E. The above lemma provides two interesting insights on $V_{i}(\vec{t})$. On one hand, the upper bound implies the information may not spread to the whole quantum system when $t\leq\mathcal{O}(1)$. On the other hand, the lower bound implies that the constant-time Hamiltonian dynamics spreads information faster than the constant-depth quantum circuit. Furthermore, we note that $V_{i}(\vec{t})$ can be efficiently computed by a polynomial classical algorithm.
Lemma 5.
There exists a classical algorithm that can exactly output $V_{i}(\vec{t})$ in $\mathcal{O}((e^{\pi teK\mathfrak{d}}/\epsilon)^{e^{\pi teK\mathfrak{d}}})$ running time such that $\|V_{i}(\vec{t})-U_{i}(\vec{t})\|\leq\epsilon$.
The proof details refer to Appendix G.
D.2 Compute $F(x)$
Definition 10 ($2$-coloring of $2$-dimensional lattice with distance $r$).
Consider a graph representing a $2$-dimensional lattice, where each vertex is assigned a color, and the entire lattice is divided into many small regions with different colors. A $2$-coloring of $2$-dimensional lattice with distance $r$ satisfies the following properties:
  • There are $2$ colors in total;
  • The distance between two regions with the same color is at least $r$.
Suppose the $n$-qubit $2$-dimensional lattice has been colored by $2$ regions, denoted by $R_{1},R_{2}$, then each region can be divided into
$R_{j}=\cup_{l=1}^{\sqrt{n}/r}R_{j}(l),$ (40)
where each zone $R_{j}(l)$ contains $r\sqrt{n}$ qubits. It is easy to check $R_{j}$ contains $S=[\sqrt{n}/r]$ zones separated by distance at least $\geq r$.
Let the zone distance $r=2M$ (given in Def. 10 and lemma 4), then for any region index $j\in[2]$ and $l\in[S]$, we have
$\left|R_{j}(l)\right|=2M\sqrt{n}.$ (41)
Recall that $\hat{\mu}(t)=\langle 0^{n}|V_{1}(t)V_{2}(t)\cdots V_{n}(t)|0^{n}\rangle$, then we can assign operators $\{V_{i}(t)\}_{i=1}^{n}$ into groups $R_{1},R_{2}$ marked by different colors, where each group is denoted by
$V(R_{j})=\bigotimes\limits_{i\in R_{j}}V_{i}(t)=\bigotimes\limits_{l=1}^{S}\left(\bigotimes_{i\in R_{j}(l)}V_{i}(t)\right)=\bigotimes\limits_{l=1}^{S}V_{R_{j}(l)}(t),$ (42)
where $V_{R_{j}(l)}(t)$ contains all $V_{i}(t)$ if the qubit index $i$ belongs to the zone $R_{j}(l)$. To deeply understand whether $V(R_{j})$ can be classically simulated, we need to evaluate the support of $V_{R_{j}(l)}(t)$.
Lemma 6.
Given the operator $V_{R_{j}(l)}(t)$ defined as Eq. 42, we have
${\rm supp}\left(V_{R_{j}(l)}(t)\right)\leq 4M\sqrt{n},$ (43)
where $M=\mathcal{O}\left(e^{\pi teK\mathfrak{d}}\log(2n/\epsilon)\right)$, meanwhile
${\rm supp}\left(V_{R_{j}(l)}(t)\right)\cap{\rm supp}\left(V_{R_{j}(q)}(t)\right)=\emptyset$ (44)
for all indexes $l\neq q\in[S]$.
The above property enables us to decouple $V(R_{j})$ into a series of operators $V_{R_{j}(l)}(t)$ whose support region does not have the overlap, and this naturally provides a classical method in simulating $\langle x|V_{R_{j}(l)}(t)|0^{n}\rangle$.
Lemma 7.
Given the operator $V_{R_{j}(l)}(t)$ defined as Eq. 42, for any for $x\in\{0,1\}^{n}$, there exists a classical algorithm that can deterministically output $\langle x|V_{R_{j}(l)}(t)|0^{n}\rangle$ within $C(n)\leq\mathcal{O}\left(\sqrt{n}2^{4M^{2}}\right)$ running time.
We leave proofs of Lemmas 6 and 7 in Appendix H.
D.3 Algorithm Complexity Analysis
Given above results, we can study the running time meanwhile proving Theorem 1. Given an observable $O=O_{1}\otimes\cdots\otimes O_{L}$ with locality $L\in[n]$, according to the algorithm given by Alg. 1, we first need to compute $V_{1}(\vec{t}),V_{2}(\vec{t}),\cdots,V_{L}(\vec{t})$. Lemma 5 implies that each $V_{i}(\vec{t})$ can be exactly computed in $\mathcal{O}((e^{\pi teK\mathfrak{d}}/\epsilon)^{e^{\pi teK\mathfrak{d}}})$ running time, where $\epsilon$ characterizes $\|U_{i}(\vec{t})-V_{i}(\vec{t})\|\leq\epsilon$. Now assigning $\epsilon\leftarrow\epsilon/2L$ and using the quantum operator synthesis inequality [101], we have
$\left\|V_{1}(\vec{t})V_{2}(\vec{t})\cdots V_{L}(\vec{t})-U_{1}(\vec{t})U_{2}(\vec{t})\cdots U_{L}(\vec{t})\right\|\leq\sum\limits_{i=1}^{L}\|V_{i}(\vec{t})-U_{i}(\vec{t})\|\leq\epsilon/2,$ (45)
which implies $\left|\hat{\mu}(\vec{t})-\mu(\vec{t})\right|\leq\mathcal{O}\left(\epsilon/2\right)$. This step requires $\mathcal{O}(L(2Le^{\pi teK\mathfrak{d}}/\epsilon)^{e^{\pi teK\mathfrak{d}}})$ running time.
After reconstructing the quantum circuit by $V_{i}(\vec{t})$, the second step is to utilize the classical Monte Carlo method to estimate $\hat{\mu}(\vec{t})$. It is shown that $\mathcal{O}(4/\epsilon^{2})$ samples from $p(x)=\left|\langle 0^{n}|V(R_{1})|x\rangle\right|^{2}$ suffice to provide a $\epsilon/2$-estimation to $\hat{\mu}(\vec{t})$, where each sample $x$ introduces
$\frac{\sqrt{n}}{4M}C(n)\leq\mathcal{O}\left(\frac{n}{4e^{\pi teK\mathfrak{d}}\log(2L/\epsilon)}2^{e^{2\pi teK\mathfrak{d}}\log^{2}(2L/\epsilon)}\right)$
computational complexity (Lemma 7). Combine all together, the classical algorithm requires
$\begin{split}&\mathcal{O}\left(\frac{4}{\epsilon^{2}}\frac{n}{4e^{\pi tK\mathfrak{d}}\log(2L/\epsilon)}\left(\frac{2L}{\epsilon}\right)^{e^{2\pi teK\mathfrak{d}}\log(2L/\epsilon)}+L\left(\frac{2Le^{\pi teK\mathfrak{d}}}{\epsilon}\right)^{e^{\pi teK\mathfrak{d}}}\right)\\ \leq&\mathcal{O}\left(\frac{n}{\epsilon^{2}}\left(\frac{2L}{\epsilon}\right)^{e^{2\pi teK\mathfrak{d}}\log(2L/\epsilon)}\right)\end{split}$ (46)
which concludes the proof.
Appendix E Proof of Lemma 4
The following lemma proves the convergence of the $K$-step cluster expansion and the support of $V_{i}(\vec{t})$ for constant times.
Theorem 5 (Formal version of Lemma 4).
Given a single qubit observable $O_{i}$, then for any $K$-step quantum dynamics driven by $\{H^{(1)},\cdots,H^{(K)}\}$ and corresponding constant time parameters $\vec{t}=\{t_{1},\cdots,t_{K}\}$, the operator $U_{i}(\vec{t})=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}}$ can be approximated by $V_{i}(\vec{t})$ such that $\|U_{i}(\vec{t})-V_{i}(\vec{t})\|\leq\epsilon\|O_{i}\|$, where
$V_{i}(\vec{t})=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}^{M}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]\right].$ (47)
Superficially, if the evolution time $t<1/(2eK\mathfrak{d})$, the number of involved cluster terms
$M=\frac{\log(1/\epsilon)-K\log(1-2teK\mathfrak{d})}{K\log(1/(2teK\mathfrak{d}))},$ (48)
otherwise we have
$M=e^{\pi teK\mathfrak{d}/\kappa}\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right],$ (49)
where the parameter $\kappa\in\mathcal{O}(1)$. Furthermore we have
$\Omega(M)\leq{\rm supp}(V_{i}(\vec{t}))\leq\mathcal{O}(4M^{2}).$ (50)
E.1 Short time Hamiltonian dynamics
Let $t=\max_{k\in[K]}\{t_{k}\}$, we first consider the scenario $\left|t\right|\leq 1/(2eK\mathfrak{d})$, where the constant $\mathfrak{d}$ represents the maximum degree of the Hamiltonian interaction graph. ($K\mathfrak{d}$ represents the maximum degree of the $m$-sized graph $G^{K}_{\bm{V}_{1}\cdots\bm{V}_{K}}$ which connects to $O_{i}$.) Now we study the convergence of the cluster expansion
$U_{i}(\vec{t})=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots\left[h_{V_{\sigma_{1}(m_{1})}}\cdots\left[h_{V_{\sigma_{K}(m_{K})}},O_{i}\right]\right]\right]$
up to index $m_{1},m_{2},\cdots m_{K}\leq M$. Specifically, let $m=m_{1}+\cdots+m_{K}$, we have
$\begin{split}\epsilon_{M}(\vec{t})=&\|U_{i}(\vec{t})-V_{i}(\vec{t})\|\\ =&\left\|\sum\limits_{\begin{subarray}{c}m_{1}\geq M+1\\ \cdots\\ m_{K}\geq M+1\end{subarray}}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]\right]\right\|\\ \leq&\sum\limits_{m_{1},\cdots,m_{K}\geq M+1}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{{\bm{\lambda}}^{\bm{V}_{1}}\cdots{\bm{\lambda}}^{\bm{V}_{K}}(2t_{1})^{m_{1}}\cdots(2t_{K})^{m_{K}}}{(\bm{V}_{1}!\cdots\bm{V}_{K}!)}\left\|O_{i}\right\|\\ \leq&\sum\limits_{m_{1},\cdots,m_{K}\geq M+1}(2t_{1})^{m_{1}}\cdots(2t_{K})^{m_{K}}\left|\mathcal{G}_{m}^{K,O_{i}}\right|\|O_{i}\|\\ \leq&\|O_{i}\|\sum\limits_{m_{1},\cdots,m_{K}\geq M+1}(2t_{1})^{m_{1}}\cdots(2t_{K})^{m_{K}}\left|eK\mathfrak{d}\right|^{m_{1}+\cdots m_{K}}\\ \leq&\|O_{i}\|\left[\sum\limits_{l\geq M+1}(2teK\mathfrak{d})^{l}\right]^{K},\end{split}$ (51)
where $t=\max_{k\in[K]}\{t_{k}\}$. The second line is valid since $\left\|\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}\right]\right\|\leq 2^{m_{1}+\cdots+m_{K}}\max\|h_{i}\|\|O_{i}\|\leq 2^{m}\|O_{i}\|_{2}$, and the fifth line comes from $\left|\mathcal{G}_{m}^{K,O_{i}}\right|\leq(eK\mathfrak{d})^{m}$ (refers to proposition 3.6 in Ref. [47]).
As a result, when $t\leq 1/(2eK\mathfrak{d})$, we have
$\epsilon_{M}(\vec{t})\leq\|O_{i}\|\frac{(2teK\mathfrak{d})^{K(M+1)}}{(1-2teK\mathfrak{d})^{K}}.$ (52)
Let $\epsilon=\frac{(2teK\mathfrak{d})^{K(M+1)}}{(1-2teK\mathfrak{d})^{K}}$, this results in
$M=\frac{\log(1/\epsilon)-K\log(1-2teK\mathfrak{d})}{K\log(1/(2teK\mathfrak{d}))}.$ (53)
E.2 Arbitrary constant time Hamiltonian dynamics
Noting that above process can be further generalized to an arbitrary constant time $t$ by means of analytic continuation. Consider the radius of a disk $R>1$, the analytic continuation can be achieved by using the map $t\mapsto t\phi(z)$, where the complex function
$\phi(z)=\frac{\log(1-z/R^{\prime})}{\log(1-1/R^{\prime})}$
maps a disk onto an elongated region along the real axis [35]. Here, the parameter $R^{\prime}>R$, and $\phi(z)$ is analytic on the closed desk $D_{R}=\{z\in\mathbb{C}:\left|z\right|\leq R\}$. Meanwhile, $\phi(z)$ satisfies $\phi(0)=0$, $\phi(1)=1$ and we select the branch ${\rm Im}(\phi(z))\leq-\pi/(2\log(1-1/R^{\prime}))$.
We consider the function
$f(z)=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}\phi(z)}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}\phi(z)}$ (54)
on the region $\left|z\right|\leq sR$ where $s\in(0,1)$. Consider a curve ${\mathcal{C}}^{\prime}=\{\left|w\right|=R\}$, according to the Cauchy integral method, we have
[…]
As a result, the truncated error can be upper bounded by
[…]
We require the following result to evaluate the upper bound of $\|f(w)\|_{2}$.
Definition 11 (Multi-variable complex analytic function).
Suppose $g:D\mapsto\mathbb{C}$ be a function on the domain $D\subset\mathbb{C}^{K}$, if for any vector $\beta\in D$, there exists a $r$-radius cylinder $P_{K}(\beta,r)$ centered on $\beta$, such that
$g(z)=\sum\limits_{\alpha_{1},\cdots,\alpha_{K}\geq 0}c_{\vec{\alpha}}(z_{1}-\beta_{1})^{\alpha_{1}}\cdots(z_{K}-\beta_{K})^{\alpha_{K}},$ (57)
then $g$ is analytic on the point $\beta=(\beta_{1},\cdots,\beta_{K})$.
Lemma 8.
Given complex values $\vec{w}=(w_{1},\cdots,w_{K})\in\mathbb{C}^{K}$, if ${\rm Im}(w_{k})\leq 1/(2eK\mathfrak{d})$ for all $k\in[K]$, we have
$\|U_{i}(\vec{w})\|\leq\frac{\|O_{i}\|}{(1-2\left|\max_{k}{\rm Im}(w_{k})\right|eK\mathfrak{d})^{K}},$ (58)
where $\mathfrak{d}$ represents the maximum degree of the interaction graph induced by Hamiltonian $H$.
Proof.
Eq. 51 provides an approximation to $U_{i}(\vec{t})$ when $\max_{k}\left|t_{k}\right|\leq 1/(2eK\mathfrak{d})$, in other word, $U_{i}(\vec{t})$ remains analytic for all complex values $t_{k}\in\mathbb{C}$ in the range $\left|t_{k}\right|<1/(2eK\mathfrak{d})$. Specifically, given any $\beta_{1},\beta_{2},\cdots,\beta_{K}\in\mathbb{R}$, we may write $U_{i}(\vec{t})=\prod_{k=1}^{K}e^{iH^{(k)}(t_{k}-\beta_{k})}e^{iH^{(k)}\beta_{k}}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}(t_{k}-\beta_{k})}e^{-iH^{(k)}\beta_{k}}$. Equivalently, we have
$U_{i}(\vec{t})=\sum\limits_{l_{1}\cdots,l_{K}\geq 0}u_{l_{1},\cdots,l_{K}}(t_{1}-\beta_{1})^{l_{1}}\cdots(t_{K}-\beta_{K})^{l_{K}}$ (59)
for some operators $u_{l_{1},\cdots,l_{K}}$, which naturally implies $U_{i}(\vec{t})$ is analytic for all complex values of $\vec{t}$ on a disk in the complex plane of radius $1/(2eK\mathfrak{d})$ around any point on the real axis.
For $\vec{w}=(w_{1},\cdots,w_{K})\in\mathbb{C}^{K}$, noting that $e^{-i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}e^{i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}=I$, then for any matrix $A$, the matrix
$e^{-i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}Ae^{i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}$
is similar to $A$, and they thus share the same spectrum information. This further implies
$\|e^{-i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}UAU^{\dagger}e^{i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}\|=\|e^{-i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}Ae^{i(w_{k}-{\rm Re}(w_{k}))H^{(k)}}\|$ (60)
for any unitary matrix $U$ (operator norm in invariant under unitary transformation). Using this property, for any $w\in\mathbb{C}^{K}$, we have
$\begin{split}\|U_{i}(\vec{w})\|=&\Big{\|}e^{-i(w_{K}-{\rm Re}(w_{K}))H^{(K)}}e^{-i{\rm Re}(w_{K})H^{(K)}}\cdots e^{-i(w_{1}-{\rm Re}(w_{1}))H^{(1)}}e^{-i{\rm Re}(w_{1})H^{(1)}}O_{i}\\ &e^{i{\rm Re}(w_{1})H^{(1)}}e^{i(w_{1}-{\rm Re}(w_{1}))H^{(1)}}\cdots e^{i{\rm Re}(w_{K})H^{(K)}}e^{i(w_{K}-{\rm Re}(w_{K}))H^{(K)}}\Big{\|}\\ =&\Big{\|}e^{-i(w_{K}-{\rm Re}(w_{K}))H^{(K)}}\cdots e^{-i(w_{1}-{\rm Re}(w_{1}))H^{(1)}}O_{i}e^{i(w_{1}-{\rm Re}(w_{1}))H^{(1)}}\cdots e^{i(w_{K}-{\rm Re}(w_{K}))H^{(K)}}\Big{\|}.\end{split}$ (61)
For square matrices $A$ and $B$, the BCH expansion enables us to write the cluster expansion to $e^{tA}Be^{-tA}$ [47] for $t\in\mathbb{R}$. As a result, we have
$\begin{split}\|U_{i}(\vec{w})\|=&\left\|\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-i(w_{k}-{\rm Re}(w_{k})))^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{K}(m_{K})}},O_{i}]\right]\right\|\\ \leq&\|O_{i}\|\sum\limits_{m_{1},\cdots,m_{K}\geq 0}\left|(2(w_{1}-{\rm Re}(w_{1})))^{m_{1}}\cdots(2(w_{K}-{\rm Re}(w_{K})))^{m_{K}}\right|\left|eK\mathfrak{d}\right|^{m_{1}+\cdots m_{K}}\\ =&\|O_{i}\|\sum\limits_{m_{1},\cdots,m_{K}\geq 0}\left|(2({\rm Im}(w_{1})))^{m_{1}}\cdots(2({\rm Im}(w_{K})))^{m_{K}}\right|\left|eK\mathfrak{d}\right|^{m_{1}+\cdots m_{K}}\\ =&\frac{\|O_{i}\|}{(1-2\left|\max_{k}{\rm Im}(w_{k})\right|eK\mathfrak{d})^{K}}.\end{split}$ (62)
Recall that
$f(w)=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}\phi(w)}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}\phi(w)}$
where $\vec{t}\in\mathbb{R}^{K}$ and ${\rm Im}(\phi(w))\leq-\pi/(2\log(1-1/R^{\prime}))$. Assign $\vec{t}\phi(w)$ to $\vec{w}$ given in Lemma 8, then Lemma 8 implies
$\begin{split}\|f(w)\|=\|U_{i}(\phi(w)\vec{t})\|&\leq\frac{\|O_{i}\|}{(1-2\left|\max_{k}{\rm Im}(t_{k}\phi(w))\right|eK\mathfrak{d})^{K}}\\ &\leq\frac{\|O_{i}\|}{(1+\pi teK\mathfrak{d}/(\log(1-1/R^{\prime})))^{K}}\end{split}$ (63)
for all $w\in C^{\prime}=\{\left|w\right|=R\}$. This further results in
[…]
where the last line follow from the fact that $\left|w-z\right|\geq R(1-s)$, $\left|z\right|\leq sR$ and $\|w\|=R$. Combine inequalities 63 and 64, we have
$\left\|f(z)-\sum\limits_{k=0}^{M}\frac{f^{(k)}(0)}{k!}z^{k}\right\|\leq\frac{\|O_{i}\|s^{M+1}}{(1+\pi teK\mathfrak{d}/(\log(1-1/R^{\prime})))^{K}(1-s)}.$ (65)
Let $\kappa=\frac{-\pi teK\mathfrak{d}}{\log(1-1/R^{\prime})}$, $R^{\prime}$ can be further expressed by
$\frac{1}{R^{\prime}}=1-e^{-\pi teK\mathfrak{d}/\kappa}.$ (66)
Since the parameter $R^{\prime}>R$, we can always select $R$ such that $(R^{\prime})^{M}(R^{\prime}-1)=2R^{M}(R-1)$ holds. Substitute this relationship into the approximation upper bound given by 65 and assign $s=1/R$, we finally obtain
$\epsilon=\frac{s^{M+1}}{\left(1+\frac{\pi teK\mathfrak{d}}{\log(1-1/R^{\prime})}\right)^{K}(1-s)}=\frac{1}{(1-\kappa)^{K}}\left(1-e^{-\pi teK\mathfrak{d}/\kappa}\right)^{M}\left(e^{\pi teK\mathfrak{d}/\kappa}-1\right).$ (67)
This implies truncating at order
$M=\frac{\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right]}{\log\left[e^{\pi teK\mathfrak{d}/\kappa}/\left(e^{\pi teK\mathfrak{d}/\kappa}-1\right)\right]}\approx e^{\pi teK\mathfrak{d}/\kappa}\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right].$ (68)
E.3 Evaluate the support size
Finally, we evaluate the support of $V_{i}(\vec{t})$. From the above sections, we observe that the support of $V_{i}$ is determined by connected cluster graphs $\mathcal{G}_{m}^{K}$. Noting that the cluster graph $\mathcal{G}_{m}^{K}$ should directly connect to the operator $O_{i}$, otherwise the commutator net may vanish. To represent such constraint, we denote $\mathcal{G}_{m}^{K}$ as $\mathcal{G}_{m}^{K,O_{i}}$, and we have
${\rm supp}(V_{i}(\vec{t}))=\bigcup\limits_{m=0}^{M}{\rm supp}(\mathcal{G}_{m}^{K,O_{i}})={\rm supp}(\mathcal{G}_{M}^{K,O_{i}}),$ (69)
where the second equality holds since $\mathcal{G}_{m}^{K,O_{i}}\subset\mathcal{G}_{m+1}^{K,O_{i}}$. A simple method for evaluating ${\rm supp}(\mathcal{G}_{M}^{K,O_{i}})$ is to measure the qubit area covered by $\mathcal{G}_{M}^{K,O_{i}}$. Recall that $\mathcal{G}_{M}^{K,O_{i}}$ contains all connected clusters (subgraphs) with size $M$, we can take $O_{i}$ as a center, and plot a circle $C_{R}$ with radius $R=M$. Then $C_{R}$ contains all connected clusters with size $\leq M$, which naturally implies the relationship
${\rm supp}(\mathcal{G}_{M}^{K,O_{i}})\leq\mathcal{O}(4M^{2}).$ (70)
On other hand, a similar idea can also be utilized to lower bound ${\rm supp}(\mathcal{G}_{M}^{K,O_{i}})$. Given the two-dimensional Hamiltonian $H=\sum_{X\in S}\lambda_{X}h_{X}$, without loss of generality, we have $\cap_{X\in S}X\neq\emptyset$, otherwise the Hamiltonian dynamics can be decoupled into two non-interactive subsystems. As a consequence, if $M\leq n$ then $\mathcal{G}_{M}^{K,O_{i}}$ contains at least one $M$-length path which corresponds to $h_{X_{1}},\cdots,h_{X_{M}}$. This provides the lower bound
${\rm supp}(\mathcal{G}_{M}^{K,O_{i}})\geq\Omega(M).$ (71)
Appendix F Circuit Depth Complexity Lower Bound
Corollary 4.
Given $K$ Hamiltonians $\{H^{(1)},\cdots,H^{(K)}\}$ defined on a two-dimensional lattice with $n$ qubits and a time series $\vec{t}=(t_{1},\cdots,t_{K})$, we denote $U(\vec{t})=\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}}$. Then the quantum circuit depth complexity of $U(\vec{t})$ satisfies
$C(U(\vec{t}))\geq\Omega\left(\pi eK\mathfrak{d}t/\kappa\right),$ (72)
where the parameters $t=\max\{t_{k}\}_{k=1}^{K}$, $\mathfrak{d}$ represents the maximum degree of the related interaction graphs and $\kappa\in\mathcal{O}(1)$.
Proof of Corollary 1.
Suppose $U(\vec{t})=\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}}$ and arbitrarily select a single qubit Pauli operator $P_{i}\in\{X,Y,Z\}$. Denoting $V_{i}(\vec{t})$ represents the truncated cluster expansion of $U(\vec{t})P_{i}U^{\dagger}(\vec{t})$ up to $M$ order (Eq. 38), according to the support lower bound on ${\rm supp}(V_{i}(\vec{t}))$ (given by lemma 4), we know that
${\rm supp}(V_{i}(\vec{t}))\geq\Omega(M)=\Omega\left(e^{\pi teK\mathfrak{d}/\kappa}\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right]\right).$ (73)
On other hand, $V_{i}(\vec{t})$ is essentially an approximation to $U(\vec{t})P_{i}U^{\dagger}(\vec{t})$ which discards all clusters with size $\geq M$. This naturally implies
${\rm supp}(V_{i}(\vec{t}))\leq{\rm supp}(U(\vec{t})P_{i}U^{\dagger}(\vec{t}))\leq 2^{C(U(\vec{t}))},$ (74)
where $C(U(\vec{t}))$ represents the quantum circuit depth complexity of $U(\vec{t})$. Finally, we have
$C(U(\vec{t}))\geq\Omega\left(\log\left(e^{\pi teK\mathfrak{d}/\kappa}\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right)\right]\right)\geq\Omega\left(\pi eK\mathfrak{d}t/\kappa\right).$ (75)
Appendix G Proof of Lemma 5
Lemma 9 (Lemma 5 in Appendix D).
There exists a classical algorithm that can exactly output $V_{i}(\vec{t})$ in $\mathcal{O}((e^{\pi teK\mathfrak{d}}/\epsilon)^{e^{\pi teK\mathfrak{d}}})$ running time such that $\|V_{i}(\vec{t})-U_{i}(\vec{t})\|\leq\epsilon$.
Proof.
Here we provide details on computing $V_{i}(\vec{t})$. We consider the polynomial expression of the function
$\begin{split}f(z)&=\prod_{k=1}^{K}e^{iH^{(k)}t_{k}\phi(z)}O_{i}\prod_{k=1}^{K}e^{-iH^{(k)}t_{k}\phi(z)}\\ &=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{K}\geq 0\end{subarray}}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{K}\in\mathcal{G}_{m}^{K,O_{i}}}\frac{\prod_{k=1}^{K}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k}\phi(z))^{m_{k}})}{\prod_{k=1}^{K}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{K}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots\left[h_{V_{\sigma_{1}(m_{1})}}\cdots\left[h_{V_{\sigma_{K}(m_{K})}},O_{i}\right]\right]\right]\\ &=\sum\limits_{l_{1}=0,\cdots,l_{K}=0}^{+\infty}A_{l_{1},\cdots,l_{K}}t_{1}^{l_{1}}\cdots t_{K}^{l_{K}}[\phi(z)]^{l_{1}+\cdots+l_{K}}\end{split}$ (76)
where the first equality comes from Eq. 36 combined with discarding disconnected clusters, and $A_{l_{1},\cdots,l_{K}}$ represents the operator which is independent to variables $\{t_{k},\phi(z)\}_{k=1}^{K}$. In Appendix E, we have known that $V_{i}(\vec{t})$ is essentially the approximation to $f(1)$ up to $M$ degree, that is
$V_{i}(\vec{t})=\sum\limits_{m=0}^{M}\frac{f^{(m)}(0)}{m!}.$ (77)
As a result, computing gradient functions $f^{(m)}(0)$ for index $m\in[M]$ suffice to exactly compute $V_{i}(\vec{t})$.
Recall that $\phi(z)=\log((1-z/R^{\prime})/(1-1/R^{\prime}))=\sum_{l=0}^{+\infty}\phi_{l}z^{l}$, where $\phi_{l}=\frac{1}{l(R^{\prime})^{l}\log(1-1/R^{\prime})}$ for $l\geq 1$ and $\phi_{0}=0$. This enables us to compute
$\begin{split}&\frac{{\rm d}^{m}f(z)}{{\rm d}z^{m}}\big{|}_{z=0}\\ =&\sum\limits_{l_{1},\cdots,l_{K}\geq 0}A_{l_{1}\cdots l_{K}}t_{1}^{l_{1}}\cdots t_{K}^{l_{K}}\frac{{\rm d}^{m}}{{\rm d}z^{m}}\left[\sum\limits_{s=0}^{\infty}\phi_{s}z^{s}\right]^{l_{1}+\cdots l_{K}}\big{|}_{z=0}\\ =&\sum\limits_{l_{1},\cdots,l_{K}\geq 0}A_{l_{1}\cdots l_{K}}t_{1}^{l_{1}}\cdots t_{K}^{l_{K}}\left[\sum\limits_{s_{1},\cdots,s_{l}\geq 1}^{\infty}\phi_{s_{1}}\phi_{s_{2}}\cdots\phi_{s_{l}}(s_{1}+\cdots+s_{l})\cdots(s_{1}+\cdots+s_{l}-m+1)z^{s_{1}+\cdots+s_{l}-m}\right]\big{|}_{z=0}\\ =&\sum\limits_{l_{1}+\cdots+l_{K}=1}^{m}A_{l_{1},\cdots,l_{K}}t_{1}^{l_{1}}\cdots t_{K}^{l_{K}}\sum\limits_{\begin{subarray}{c}s_{1},\cdots,s_{l}\geq 1\\ s_{1}+\cdots+s_{l}=m\end{subarray}}\phi_{s_{1}}\cdots\phi_{s_{l}}m!,\end{split}$ (78)
where the index $l=l_{1}+\cdots+l_{K}$. Now let us explain the third equality. When $z=0$, only terms with $s_{1}+\cdots+s_{l}=m$ may not vanish. Meanwhile $l>m$ may result in some non-negative index $s_{l^{*}}=0$ for $l^{*}\in[l]$ which implies $\phi_{s_{l^{*}}}=0$, as a consequence, we have $l\leq m$. Noting that the nested commutator $A_{l_{1},\cdots,l_{K}}t_{1}^{l_{1}}\cdots t_{K}^{l_{K}}$ can be numerically evaluated in time $e^{\mathcal{O}(m)}$. We refer readers to Appendix A in Ref. [35] to find more details on computing $A_{l_{1},\cdots,l_{K}}$. The derivative function $\frac{1}{M!}\frac{{\rm d}^{M}f(z)}{{\rm d}z^{M}}\big{|}_{z=0}$ thus can be exactly computed in $\mathcal{O}\left(\exp(M)\right)$ classical running time, with
$M=e^{\pi teK\mathfrak{d}/\kappa}\log\left[\frac{1}{\epsilon}\frac{e^{\pi teK\mathfrak{d}/\kappa}-1}{(1-\kappa)^{K}}\right].$
Appendix H Details on computing $F(x)$
H.1 Proof of Lemma 6
Lemma 10 (Lemma 6 in the main file).
Given the operator $V_{R_{j}(l)}(t)$ defined as Eq. 42, we have
${\rm supp}\left(V_{R_{j}(l)}(t)\right)\leq 4M\sqrt{n},$ (79)
where $M=\mathcal{O}\left(e^{\pi teK\mathfrak{d}}\log(2n/\epsilon)\right)$, meanwhile
${\rm supp}\left(V_{R_{j}(l)}(t)\right)\cap{\rm supp}\left(V_{R_{j}(q)}(t)\right)=\emptyset$ (80)
for all indexes $l\neq q\in[S]$.
Proof.
In the two-dimensional lattice, the proposed coloring method enables the region $R_{j}(l)$ being a ($\sqrt{n}\times 2M$) rectangle area, where each qubit relates to an operator $V_{i}(t)$. According to the lemma 4, it is shown that the support of $V_{i}(t)$ is upper bounded by $\mathcal{O}(M^{2})$. As a result, most of ${\rm supp}(V_{i}(t))$ have the overlap and can thus be contracted. Finally, ${\rm supp}\left(V_{R_{j}(l)}(t)\right)$ is only characterized by operators $\{V_{i}(t),i\in\partial R_{j}(l)\}$ combined with $R_{j}(l)$, that is
${\rm supp}\left(V_{R_{j}(l)}(t)\right)=\bigcup\limits_{i\in\partial R_{j}(l)}{\rm supp}\left(V_{i}(t)\right)+\left|R_{j}(l)\right|\leq 4M\sqrt{n},$ (81)
where $\partial R_{j}(l)$ is the boundary of the region $R_{j}(l)$. The visualization of the above statement is provided by Fig. 1.
Recall that ${\rm supp}\left(V_{R_{j}(l)}(t)\right)$ is essentially a quasi-1D-rectangular in the two-dimensional lattice, then the short-side length of ${\rm supp}\left(V_{R_{j}(l)}(t)\right)$ is upper bounded by $\leq 4M$. Furthermore, from Def. 10, we know that
${\rm dist}\left(R_{j}(l),R_{j}(q)\right)\geq r=2M,$ (82)
which naturally implies
${\rm supp}\left(V_{R_{j}(l)}(t)\right)\cap{\rm supp}\left(V_{R_{j}(q)}(t)\right)=\emptyset.$
for all $l\neq q$. ∎
H.2 Proof of lemma 7
Proof.
We first label all qubits contained in the region $V_{R_{j}(l)}$ by $(q_{0},\cdots,q_{4M\sqrt{n}-1})$ using the row major order. Then we consider a $M\times 4M$ window $W$ swiping along the row index. According to the result given in Lemma 4, it is shown that $\left|{\rm supp}\left(V_{i}(t)\right)\right|\leq 4M^{2}$. At the initial stage, suppose the window $W$ only covers qubit set $\mathcal{W}=\{q_{0},\cdots,q_{4M^{2}-1}\}$, then the support size of $V_{i}(t)$ implies if $q_{i}\notin\mathcal{W}$, then $V_{q_{i}}(t)$ may not affect the measurement result of $l_{0}=\{q_{0},\cdots q_{4M-1}\}$ which represents the first row within the window $W$. This further implies that computing $\langle x_{l_{0}}|V_{l_{0}}(t)|0_{l_{0}}\rangle$ can be fixed into a small subspace, and the state vector simulator has runtime approximately $\mathcal{O}(2^{4M^{2}})$ per each gate or one-qubit measurement [36]. Swiping the window $W$ from $l_{0}$ to $l_{\sqrt{n}-1}$, $\langle x|V_{R_{j}(l)}(t)|0^{n}\rangle$ can be deterministically computed by a classical algorithm with $\tilde{\mathcal{O}}\left(\sqrt{n}2^{4M^{2}}\right)$ running time.
Appendix I Limitations of NISQ Algorithms on current quantum devices
Here, we study the Hamiltonian simulation algorithm on near-term quantum devices in the context of a noisy environment. We suppose each quantum gate is affected by a local Pauli channel. For the sake of clarity, we begin by presenting the definitions of the local Pauli channel.
Definition 12 (Local Pauli channel).
Let $\mathcal{N}_{i}$ denote a local Pauli channel and the action of $\mathcal{N}_{i}$ is random local Pauli operators $P$ acting on the $i$-th qubit according to a specific channel parameter $\{q(P)\}$, where $P\in\{\mathbb{I},\sigma^{x},\sigma^{y},\sigma^{z}\}$. Specifically, the action of $\mathcal{N}_{i}$ is given by
$\mathcal{N}_{i}(P)=q(P)P$ (83)
for $P\in\{\mathbb{I},\sigma^{x},\sigma^{y},\sigma^{z}\}$, where $q(P)\in(-1,1)$. The noise strength in this model is represented by a single parameter $q=\max_{P\in\{\sigma^{x},\sigma^{y},\sigma^{z}\}}\left|q(P)\right|$.
Definition 13 (Quantum Circuit affected by Pauli Channel).
We assume that the noise in the quantum device is modeled by a Pauli channel $\mathcal{N}_{i}$ with strength $q$. Let $\mathcal{U}$ be a causal slice, and let $\mathcal{N}\circ\mathcal{U}=\left(\otimes_{i=1}^{n}\mathcal{N}_{i}\right)\circ\mathcal{U}$ be the representation of a noisy circuit layer. We define the $d$-depth noisy quantum state with noise strength $q$ as
$\rho_{q,d}=\mathcal{N}\circ\mathcal{U}_{d}\circ\mathcal{N}\circ\mathcal{U}_{d-1}\circ\cdots\circ\mathcal{N}\circ\mathcal{U}_{1}(|0^{n}\rangle\langle 0^{n}|).$ (84)
Quantum error mitigation is necessary due to imperfections in quantum devices to correct the bias caused by noise. The fundamental concept is to correct the impact of quantum noise through classical post-processing of measurement results, without mid-circuit measurements and adaptive gates as in standard error correction. Here, we argue that the existing error mitigation strategies might require a number of samples $\rho_{q,d}$ that scales exponentially with the number of gates in the light-cone of the observable of interest. This thus losses the original quantum advantages compared to the proposed classical simulation algorithm which only requires quasi-polynomial time. We extend previous results given by Ref [80] to a more general Pauli channel. We first review some related lemmas, then give the generalized result and main result (given in the main file) on the quantum error-mitigation overheads.
I.1 Involved Lemmas
We require following lemmas to support our proof.
Lemma 11 (Ref. [80]).
Let $P_{0},\cdots,P_{N}$ be probability measures on some state space $X$ such that
$\frac{1}{N+1}\sum\limits_{k=0}^{N}D(P_{k}\|P_{0})\leq\alpha\log(N)$ (85)
for $10<\alpha<$ 1. Then the minimum average probability of error over tests $\psi:X\mapsto\{0,1,\cdots,N\}$ that distinguish the probability distributions $P_{0},\cdots,P_{N}$ which we define as
$\bar{p}_{e,N}={\rm inf}_{\psi}\frac{1}{N+1}\sum\limits_{j=0}^{N}P_{j}(\psi\neq j)$ (86)
satisfies
$\bar{p}_{e,N}\geq\frac{\log(N+1)-\log(2)}{\log(N)}-\alpha.$ (87)
Lemma 12 (Lemma 6 in [102]).
Consider a single instanoise channel $\mathcal{N}=\mathcal{N}_{1}\otimes\cdots\otimes\mathcal{N}_{n}$ where each local noise channel $\{\mathcal{N}_{j}\}_{j=1}^{n}$ is a Pauli noise channel that satisfies $\mathcal{N}_{j}(\sigma)=q_{\sigma}\sigma$ for $\sigma\in\{X,Y,Z\}$ and $q_{\sigma}$ be the Pauli strength. Then we have
$D_{2}\left(\mathcal{N}(\rho)\|\frac{I^{\otimes n}}{2^{n}}\right)\leq q^{2c}D_{2}\left(\rho\|\frac{I^{\otimes n}}{2^{n}}\right),$ (88)
where $D_{2}(\cdot\|\cdot)$ represents the $2$-Renyi relative entropy, $q=\max_{\sigma}q_{\sigma}$ and $c=1/(2\ln 2)$.
Lemma 13.
Given an arbitrary $n$-qubit density matrix and maximally mixed state $I^{\otimes n}/2^{n}$, we have
$D\left(\rho\|I^{\otimes n}/2^{n}\right)\leq D_{2}\left(\rho\|I^{\otimes n}/2^{n}\right),$ (89)
where $D(\cdot\|\cdot)$ denotes the relative entropy and $D_{2}(\cdot\|\cdot)$ denotes the $2$-Renyi relative entropy.
Proof.
Given quantum states $\rho$ and $\sigma$, the quantum $2$-Renyi entropy
$D_{2}(\rho\|\sigma)=\log{\rm Tr}\left[\left(\sigma^{-1/4}\rho\sigma^{-1/4}\right)^{2}\right].$ (90)
When $\sigma=I^{\otimes n}/2^{n}$, we have $D_{2}(\rho\|I^{\otimes n}/2^{n})=\log{\rm Tr}\left[\left((I^{\otimes n}/2^{n})^{-1}\rho^{2}\right)\right]=n+\log{\rm Tr}[\rho^{2}]$. Noting that the function $y=x^{2}-x\log x\geq 0$ when $x\in[0,1]$, and this implies ${\rm Tr}(\rho^{2})\geq{\rm Tr}(\rho\log\rho)$. Finally, we have
$D\left(\rho\|I^{\otimes n}/2^{n}\right)=n+{\rm Tr}\left[\rho\log\rho\right]+n\leq{\rm Tr}\left[\rho^{2}\right]+n=D_{2}\left(\rho\|I^{\otimes n}/2^{n}\right).$ (91)
Lemma 14 (Ref. [103]).
Let $\epsilon\in(0,1)$ and $\delta\in(0,1)$. Suppose there exists a POVM $\{M_{\sigma}d\sigma\}$ on $(\mathbb{C}^{2^{n}})^{\otimes m}$ such that for any quantum state $\rho$,
[…]
This implies the sample complexity lower bound
$m\geq\Omega\left(\frac{2^{3n}(1-\epsilon)^{2}}{\epsilon^{2}}\right).$ (93)
I.2 Generalize the Theorem 1 in Ref [80] to Pauli channel
Fact 1 (Generalized result to Ref. [80]).
Let $\mathcal{A}$ be an error mitigation algorithm that takes as input $m$ noisy quantum state copies prepared by a $d$-depth noisy quantum circuit that affected by local Pauli noise channels with strength $q$, and a set of Hermitian observables. The error mitigation algorithm $\mathcal{A}$ requires $m\geq\Omega(q^{-2d})$ copies of noisy states in the worst-case scenario over the choice of observable sets.
The basic idea is to construct a polynomial reduction to the quantum state discrimination problem [80]. Let us consider an error mitigation problem. Given the quantum state set $F_{\rho}=\{\rho_{0},\rho_{1},\cdots,\rho_{N}\}$, where $\rho_{x}=|x\rangle\langle x|$ when $x<N$ and $\rho_{N}=I_{n}/2^{n}$ with $N=2^{n}$, as the input of a noiseless quantum circuit $C$, and utilize a set of observables $\{CZ_{i}C^{\dagger}\}_{i=1}^{n}$ to measure the output states $C(\rho_{x})$. The quantum error mitigation algorithm should output the estimation $o_{j}$ such that $\left|o_{j}-{\rm Tr}(C(\rho_{x})CZ_{j}C^{\dagger})\right|\leq\epsilon$. Now we show that a noisy state identification problem can be solved by quantum error mitigation algorithm. Consider an arbitrary $\rho_{x}\in F_{\rho}$, we may have two scenarios:
  • If the unknown quantum state $\rho_{x}$ whose index satisfies $x<N$, we have $y_{j}={\rm Tr}(C(\rho_{x})CZ_{j}C^{\dagger})=1-2x_{j}$, where $x_{j}$ represents the $j$-th bit within $x$;
  • Else $\rho_{N}=I_{n}/2^{n}$ resulting in $y_{j}={\rm Tr}(C(\rho_{N})CZ_{j}C^{\dagger})=0$.
Randomly sample a quantum state $\rho_{x}\in F_{\rho}$, we denote $\hat{y}=(y_{1},\cdots,y_{n})$ and $P_{x}(\hat{y})$ represents the probability distribution on measuring the result $\hat{y}$. As a result, if a quantum error mitigation algorithm can successfully recover every $o_{j}$ for $j\in[n]$, this enables us to uniquely identify the unknown quantum state $\rho_{x}$ from the distribution $P_{x}(\hat{y})$. Then we can utilize Fano’s lower bound for quantum state identification problem (Lemma 11). Specifically, we have
$\begin{split}\frac{1}{N+1}\sum\limits_{k=0}^{N}D(P_{k}\|P_{0})\leq&\frac{1}{N+1}\sum\limits_{k=0}^{N}D\left(\Phi^{\otimes m}_{C,q}(\rho_{k})\|(I_{n}/2^{n})^{\otimes m}\right)\\ \leq&\frac{1}{N+1}\sum\limits_{k=0}^{N}D_{2}\left(\Phi^{\otimes m}_{C,q}(\rho_{k})\|(I_{n}/2^{n})^{\otimes m}\right)\\ \leq&\frac{1}{N+1}\sum\limits_{k=0}^{N}mq^{2cd}D_{2}\left(\rho_{k}\|I_{n}/2^{n}\right)\\ =&q^{2cd}mn\\ =&q^{2cd}m\log N,\end{split}$ (94)
where $d$ represents the depth of quantum circuit $C$. Let $\alpha=q^{2cd}m$, then in order for the test to have a constant failure probability $\delta$, it takes at least $m\geq q^{-2cd}(1-\delta)$ copies.
I.3 A sample complexity lower bound related to approximation error and circuit depth
Problem 2.
Consider a pure quantum state packing net $\{\rho_{0},\cdots,\rho_{\left|\Omega\right|}\}$ such that for $\frac{1}{2}\|\rho_{i}-\rho_{j}\|\geq 2\epsilon$ for any $i\neq j$, and a $d$-depth quantum circuit $\mathcal{C}$ affected by Pauli channel $\mathcal{N}$. Suppose that a distinguisher has knowledge of $\mathcal{C}$ and $\mathcal{N}$, and is given access to copies of the quantum state $\Phi_{\mathcal{C},q}(\rho_{i})$, with some unknown index $i\in[\left|\Omega\right|]$. What is the fewest number of copies of $\Phi_{\mathcal{C},q}(\rho_{i})$ sufficing to identify $i\in[\left|\Omega\right|]$ with high probability?
Now we discuss how to utilize the quantum error mitigation algorithm to solve the above problem. Suppose the noisy state $\Phi_{\mathcal{C},q}(\rho_{i})$ is provided, we focus on its quantum mean value on observables
$\{C^{\dagger}\rho_{0}C,\cdots,C^{\dagger}\rho_{\left|\Omega\right|}C\}$
that is to estimate $\{{\rm Tr}\left(\Phi_{\mathcal{C},q}(\rho_{i})C^{\dagger}(\rho_{j})C\right)\}$ for $j\in[N]$. If a quantum error mitigation algorithm $\mathcal{A}$ can recover the quantum mean value, then we have the map
$\left\{{\rm Tr}\left(\Phi_{\mathcal{C},q}(\rho_{i})C^{\dagger}(\rho_{j})C\right)\right\}\mapsto\{{\rm Tr}\left[\rho_{i}\rho_{j}\right]\}.$ (95)
According to our assumption, all quantum states $\rho_{i}$ comes from a packing-net, then for any $i\neq j$, we have ${\rm Tr}(\rho_{i}\rho_{j})=\sqrt{1-d^{2}_{tr}(\rho_{i},\rho_{j})}\leq\sqrt{1-4\epsilon^{2}}\leq 1-2\epsilon^{2}$. Otherwise we have ${\rm Tr}(\rho_{i}\rho_{i})\geq 1-\epsilon^{2}$. As a result, a quantum error mitigation algorithm can be used to identify the index $i$ hidden in the noisy state $\Phi_{\mathcal{C},q}(\rho_{i})$, which thus can solve Problem 2. The sample complexity of Problem 2 can be used to benchmark the sample complexity lower bound of the quantum error mitigation problem.
Theorem 6.
Let $\mathcal{A}$ be an input state-agnostic error mitigation algorithm that takes as input $m$ copies noisy quantum states produced by a $d$-depth quantum circuit $\mathcal{C}$ affected by $q$-strength local Pauli noise channels, and a set of observables $\{O\}$. Suppose the algorithm $\mathcal{A}$ is able to produce estimates $\{\hat{o}\}$ such that $\left|\hat{o}-\langle o\rangle\right|\leq\epsilon$. Then the sample complexity
$m\geq\min\left\{\frac{q^{-2cd}(1-\eta)^{2}}{2n},\frac{2^{3n}(1-\epsilon)^{2}}{\epsilon^{2}}\right\}$ (96)
in the worst-case scenario over the choice of the observable set, where $c=1/(2\ln 2)$ and $\eta\in\mathcal{O}(1)$.
Proof.
Randomly select $\rho_{i}$ and $\rho_{j}$ from the $\epsilon$-packing net, we consider the sample complexity $m$ in distinguishing quantum states $\Phi_{\mathcal{C},q}(\rho_{i})$ and $\Phi_{\mathcal{C},q}(\rho_{j})$. When their trace distance is quite large, let $\eta\in(0,1)$ and we have
$\begin{split}1-\eta&\leq\frac{1}{2}\left\|\Phi_{\mathcal{C},q}(\rho_{i})^{\otimes m}-\Phi_{\mathcal{C},q}(\rho_{j})^{\otimes m}\right\|_{1}\\ &\leq\frac{1}{2}\left(\left\|\Phi_{\mathcal{C},q}(\rho_{i})^{\otimes m}-(I_{n}/2^{n})^{\otimes m}\right\|_{1}+\left\|\Phi_{\mathcal{C},q}(\rho_{j})^{\otimes m}-(I_{n}/2^{n})^{\otimes m}\right\|_{1}\right)\\ &\leq\frac{1}{\sqrt{2}}\left(D^{1/2}\left(\Phi^{\otimes m}_{\mathcal{C},q}(\rho_{i})\|(I_{n}/2^{n})^{\otimes m}\right)+D^{1/2}\left(\Phi^{\otimes m}_{\mathcal{C},q}(\rho_{i})\|(I_{n}/2^{n})^{\otimes m}\right)\right),\end{split}$ (97)
where the second line comes from the triangle inequality and the third line comes from the Pinsker’s inequality. Using Lemmas 13 and 12, we have
$1-\eta\leq\frac{1}{\sqrt{2}}\left(D^{1/2}_{2}\left(\Phi^{\otimes m}_{\mathcal{C},q}(\rho_{i})\|(I_{n}/2^{n})^{\otimes m}\right)+D^{1/2}_{2}\left(\Phi^{\otimes m}_{\mathcal{C},q}(\rho_{i})\|(I_{n}/2^{n})^{\otimes m}\right)\right)\leq\sqrt{2nm}q^{cd},$ (98)
where $d$ represents the quantum circuit depth of $\mathcal{C}$. As a result we have
$m\geq\frac{q^{-2cd}(1-\eta)^{2}}{2n}.$ (99)
On other hand, when quantum states $\Phi_{\mathcal{C},q}(\rho_{i})$ and $\Phi_{\mathcal{C},q}(\rho_{j})$ are very close, that is $\frac{1}{2}\|\Phi_{\mathcal{C},q}(\rho_{i})-\Phi_{\mathcal{C},q}(\rho_{j})\|_{1}\leq\epsilon$ (this is possible since a CPTP map reduces the trace distance), Lemma 14 implies the sample complexity
$m\geq\frac{2^{3n}(1-\epsilon)^{2}}{\epsilon^{2}}.$ (100)
Combine above inequalities together, we finally have
$m\geq\min\left\{\frac{q^{-2cd}(1-\eta)^{2}}{2n},\frac{2^{3n}(1-\epsilon)^{2}}{\epsilon^{2}}\right\}.$ (101)
Appendix J Classical Simulation for $2$D Fermi-Hubbard model
The Fermionic Hubbard model has served as a paradigmatic example for strongly correlated problems. Specifically, its Hamiltonian is given by
$H_{FH}=-\tau\sum\limits_{(i,j)\in E,\sigma\in\{\uparrow,\downarrow\}}(a_{i\sigma}^{\dagger}a_{j\sigma}+a_{j\sigma}^{\dagger}a_{i\sigma})+U\sum\limits_{i\in V}n_{i\uparrow}n_{i\downarrow},$ (102)
where $\tau,U$ are coupling parameters of the model, $n_{j\sigma}=a^{\dagger}_{j\sigma}a_{j\sigma}$, and the fermionic creation operators $a_{i\sigma}$ satisfy $a_{i\sigma}^{\dagger}a_{j\tau}+a_{j\tau}a_{i\sigma}^{\dagger}=\delta_{ij}\delta_{\sigma\tau}$.
[…]
Using the above representation, the nearest-neighbor couplings for horizontal edges maps to $5$-local operators:
$a_{k+1}^{\dagger}a_{k}+a^{\dagger}_{k}a_{k+1}\mapsto\frac{1}{2}Y_{k}^{\rightarrow}\left(Z_{k}^{\downarrow}Z_{k+1}^{\uparrow}-Z_{k}^{\uparrow}Z_{k}^{\leftarrow}Z_{k+1}^{\rightarrow}Z_{k+1}^{\downarrow}\right),$ (103)
while the vertical nearest-neighbour couplings are encoded by $7$-local operators:
$a_{j}^{\dagger}a_{k}+a_{k}^{\dagger}a_{j}\mapsto\frac{1}{2}\left(Z_{k}^{\leftarrow}Z_{k}^{\rightarrow}Z_{k}^{\uparrow}Z_{j}^{\leftarrow}Z_{j}^{\rightarrow}Z_{j}^{\downarrow}-I\right).$ (104)
Finally, the onsite interactions
$n_{i\uparrow}n_{i\downarrow}\mapsto\frac{1}{4}\left(I-Z_{k}^{\leftarrow}Z_{k}^{\uparrow}Z_{k}^{\rightarrow}Z_{k}^{\downarrow}\right)\left(I-Z_{k^{\prime}}^{\leftarrow}Z_{k^{\prime}}^{\uparrow}Z_{k^{\prime}}^{\rightarrow}Z_{k^{\prime}}^{\downarrow}\right),$ (105)
where the primed indices correspond to fermions in spin down lattice and the unprimed ones to the sites in the spin up lattice. This implies each onsite term can be represented by a $8$-local Pauli operator.
J.1 VQE Algorithm Simulation
Here, we consider to utilize the Hamiltonian variational (HV) ansatz to estimate the ground state energy of Fermi-Hubbard model. The HV ansatz is based on intuition from the quantum adiabatic theorem, which states that one can evolve from the ground state of a Hamiltonian $H_{A}$ to the ground state of another Hamiltonian $H_{B}$ by applying a sequence of evolutions of the form $e^{-itH_{A}}$ and $e^{-itH_{B}}$ for sufficiently small time $t$. In our case, the HV ansatz starts from the ground state of the non-interacting Hubbard Hamiltonian $(U=0)$ which is essentially a slater determinant quantum state. Each layer of the HV ansatz is constructed by
$e^{-iH_{v}t_{v}}e^{-iH_{h}t_{h}}e^{-iH_{o}t_{o}},$ (106)
where time series $\{t_{v},t_{h},t_{o}\}$, $H_{V}$ is the vertical hopping term, $H_{h}$ is the horizontal hopping term and $H_{o}$ is the onsite term. Suppose the HV ansatz contains $p$ layers, the initial quantum state is $|\phi\rangle$, then the VQE algorithm minimizes the energy function
$E(\vec{t})=\langle\phi|\prod\limits_{j=1}^{p}e^{iH_{v}t^{(j)}_{v}}e^{iH_{h}t^{(j)}_{h}}e^{iH_{o}t^{(j)}_{o}}H_{FH}\prod\limits_{j=1}^{p}e^{-iH_{v}t^{(j)}_{v}}e^{-iH_{h}t^{(j)}_{h}}e^{-iH_{o}t^{(j)}_{o}}|\phi\rangle$ (107)
in each optimization step. It is shown that Hamiltonian $H_{\rm FH}$ can be decomposed by linear combinations of local Pauli operators, and so the energy function is a sum of $\mathcal{O}(n^{2})$ mean values of local observable.
Corollary 5.
Given a two-dimensional Fermi-Hubbard model defined on a $(a\times b)$-sized lattice, a $p$-depth Hamiltonian Variational ansatz (given by Eq. 106) with parameters $\{t^{(j)}_{v},t^{(j)}_{h},t^{(j)}_{o}\}_{j=1}^{p}\in[-\pi,\pi]^{3p}$ and a slater determinant initial state, then each step of the corresponding VQE program can be simulated by a classical algorithm with a run time
$\mathcal{O}\left(\frac{4ab}{\epsilon^{2}}\left(\frac{2L}{\epsilon}\right)^{e^{4\pi^{2}ep\mathfrak{d}}\log(2L/\epsilon)}\right),$ (108)
where the constant $\mathfrak{d}$ represents the maximum degree of the interaction graph induced by $H_{\rm FH}$ and the locality $L\leq 8$.
Proof.
The superfast encoding method may encode a $(a\times b)$-sized Hamiltonian into a $(2a\times 2b)$-sized Hamilotnian. Then taking $t=2\pi$, $n=4ab$ into Theorem 1 may conclude the result directly. ∎
When the HV ansatz depth $p\leq\mathcal{O}(1)$, above result implies VQE algorithm can be efficiently simulated by a classical algorithm, and this further suggests VQE algorithms may loss exponential speed-up in terms of the system size.
J.2 Quantum State Property Simulation
Given a $2$-dimensional Fermi-Hubbard model, determining its quantum phase diagram under specific external parameter is of significance. Suppose the ground state $|\psi_{g}\rangle$ of $H_{\rm FH}$ has been prepared by a VQE approach, that is
$|\psi_{g}\rangle=\prod\limits_{j=1}^{p}e^{-iH_{v}t^{(j)}_{v}}e^{-iH_{h}t^{(j)}_{h}}e^{-iH_{o}t^{(j)}_{o}}|\phi\rangle.$ (109)
The ground state property can be characterized by the value of $\langle\psi_{g}|O|\psi_{g}\rangle$, where $O$ represents the target order parameter. For example, observables related to metal-insulator transition, Friedel oscillations and antiferromagnetic orders are general local [83], while observables related to the spin-charge separation, local-gapped phases and other complex topological quantum phases are general global [85, 86]. Our classical algorithm can provide an estimation to $\langle\psi_{g}|O|\psi_{g}\rangle$, in both local (symmetry breaking phase) and global (topological phase) scenarios.
Appendix K Classical Simulation for QAOA
In theoretical computational science, constraint satisfaction problems encompass a wide range of typical problems, such as Maximum Cut, Maximum Independent Set, and Graph Coloring [105]. These problems define their constraints as clauses, with a candidate solution represented by a specific assignment of the corresponding binary variables. The objective of these problems is to find an optimal assignment that maximizes the number of satisfied clauses. In other words, solving a constraint satisfaction problem can be reformulated as optimizing a quadratic function involving binary variables. However, finding the exact solution is widely recognized as an $\rm NP$-hard problem [106]. Consequently, an alternative approach is to seek an approximate solution. Inspired by the quantum annealing process [107], QAOA was proposed and applied to solve constraint satisfaction problems. Although the prospects of achieving quantum advantages through QAOA remain unclear, it provides a simple paradigm for optimization that can be implemented on near-term quantum devices.
Here, we focus on the MaxCut problem.
Definition 14 (Maximum Cut problem).
Considering an unweighted $\mathfrak{d}$-regular graph $G=(V,E)$ with the vertices set $V=\{v_{1},\cdots,v_{n}\}$ and the edges set $E=\{e_{i,j}\}$, the Maximum Cut problem aims at dividing all vertices into two disjoint sets such that maximizing the number of edges that connect the two sets. In the context of QAOA, the problem-oriented Hamiltonian $H_{A}^{\rm{MaxCut}}$ is defined as
$H_{A}^{\rm{MaxCut}}=\frac{1}{2}\sum_{e_{i,j}\in E}(\mathbb{I}^{\otimes n}-Z_{i}\otimes Z_{j}),$ (110)
and mixer $H_{B}=\sum\limits_{i=1}^{n}X_{i}$.
Subsequently, by iteratively applying $H_{A}$ and $H_{B}$ to the initial state $\rho$ for $p$ rounds, the QAOA objective function is given by the following expectation value
$f(\vec{\beta},\vec{\gamma})=\mathrm{Tr}\left[H_{A}U(\vec{\beta},\vec{\gamma})\rho U(\vec{\beta},\vec{\gamma})^{\dagger}(\bm{\theta})\right],$ (111)
where $\rho=(|+\rangle\langle+|)^{\otimes n}$ denotes the uniform superposition over computational basis states and the QAOA circuit
$U(\vec{\beta},\vec{\gamma})=\prod_{k=1}^{p}e^{-i\beta_{k}H_{A}}e^{-i\gamma_{k}H_{B}}.$ (112)
The statistical estimation of $f(\vec{\beta},\vec{\gamma})$ can be achieved by repeating the aforementioned process with identical parameters and computational basis measurements. After defining $f(\vec{\beta},\vec{\gamma})$, the next step involves iteratively updating $\vec{\beta},\vec{\gamma}$ through classical optimization methods to maximize $f(\vec{\beta},\vec{\gamma})$ and obtain the global maximum point
$(\vec{\beta},\vec{\gamma})^{*}=\arg\max_{\bm{\vec{}}{\beta},\vec{\gamma}\in\mathcal{D}}f(\vec{\beta},\vec{\gamma}),$ (113)
where the domain $\mathcal{D}=[0,2\pi]^{2p}$.
Since all Pauli terms in $H_{A}$ are local operators, it is interesting to note that such local property enables our algorithm to bypass the $2$D constraint. Specifically, one can estimate $\langle+^{n}|U^{\dagger}(\vec{\beta},\vec{\gamma})H_{A}U(\vec{\beta},\vec{\gamma})|+^{n}\rangle$ by computing
$\langle+^{n}|U^{\dagger}(\vec{\beta},\vec{\gamma})Z_{i}Z_{j}U(\vec{\beta},\vec{\gamma})|+^{n}\rangle$ (114)
for $e_{ij}\in E$. Let $\vec{t}=(\vec{\beta},\vec{\gamma})\in[0,2\pi]^{2p}$ and using Eq. 38, we have
$V_{i,j}(\vec{t})=\sum\limits_{\begin{subarray}{c}m_{1}\geq 0\\ \cdots\\ m_{2p}\geq 0\end{subarray}}^{M}\sum\limits_{\bm{V}_{1}\cdots,\bm{V}_{2p}\in\mathcal{G}_{m}^{2p,Z_{i}Z_{j}}}\frac{\prod_{k=1}^{2p}(\bm{\lambda}^{\bm{V}_{k}}(-it_{k})^{m_{k}})}{\prod_{k=1}^{2p}\bm{V}_{k}!m_{k}!}\sum\limits_{\begin{subarray}{c}\sigma_{1}\in\mathcal{P}_{m_{1}}\\ \cdots\\ \sigma_{K}\in\mathcal{P}_{m_{2p}}\end{subarray}}\left[h_{V_{\sigma_{1}(1)}},\cdots[h_{V_{\sigma_{2p}(m_{2p})}},Z_{i}Z_{j}]\right],$ (115)
where $M\leq\mathcal{O}\left(e^{2\pi ep\tau\mathfrak{d}}\log^{2}(1/\epsilon)\right)$ (according to lemma 4), with $\tau=\max\{\left|\beta_{k}\right|,\left|\gamma_{k}\right|\}_{k=1}^{p}$. Using lemma 5, a $\epsilon$-approximation to $\langle+^{n}|V_{ij}(\vec{t})|+^{n}\rangle$ can be computed in $\tilde{\mathcal{O}}((e^{2\pi e\tau p\mathfrak{d}}/\epsilon)^{e^{2\pi e\tau p\mathfrak{d}}})$ running time. Let $\epsilon$ to $\epsilon/\left|E\right|$, the $\epsilon$-approximation to the objective function $f$ can be obtained in
$\mathcal{O}((e^{2\pi e\tau p\mathfrak{d}}\left|E\right|/\epsilon)^{e^{2\pi e\tau p\mathfrak{d}}})$ (116)
classical running time.
Appendix L Dequantization on Guided Local Hamiltonian Problem
L.1 Ancilla-Free Hadamard Test
Here, we consider to design a classical algorithm in simulating the Hadamard Test algorithm when the target problem has the particle number preserved property. Specifically, we focus on a class of Hamiltonians (quantum lattice model and electronic structure model) $H$ with the property $[H,\sum_{i}n_{i}]=0$, where $n_{i}$ represents the particle number operator on the $i$-th site. Suppose the quantum system has $n$ spin orbitals, and the initial state is given by a semi-classical state $|\psi_{c}\rangle$ (used in Refs. [37, 38]) with the particle number $P\geq 0$. Our target is to estimate both real part and imaginary part of $\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle$. In general, the quantum Hadamard Test algorithm requires the controlled $e^{-iHt}$ operation, however, the particle number preserving property enables us to bypass the requirement of ancilla qubit.
In detail, let the vacuum state $|\Omega\rangle=|0^{n}\rangle$, then the particle number symmetry enables the relationship
$e^{-iHt}|\Omega\rangle=|\Omega\rangle.$ (117)
Starting from the quantum state
$|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|\Omega\rangle+|\psi_{c}\rangle\right),$ (118)
apply the operator $e^{-iHt}$ to the quantum system, then the quantum system becomes to
$e^{-iHt}|\psi_{1}\rangle=\frac{1}{\sqrt{2}}\left(|\Omega\rangle+e^{-iHt}|\psi_{c}\rangle\right).$ (119)
Finally, we have
${\rm Re}\left[\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle\right]=\langle\psi_{1}|e^{iHt}\left(|\Omega\rangle\langle\psi_{c}|+|\psi_{c}\rangle\langle\Omega|\right)e^{-iHt}|\psi_{1}\rangle,$ (120)
${\rm Im}\left[\langle\psi_{c}|e^{-iHt}|\psi_{c}\rangle\right]=i\langle\psi_{1}|e^{iHt}\left(|\psi_{c}\rangle\langle\Omega|-|\Omega\rangle\langle\psi_{c}|\right)e^{-iHt}|\psi_{1}\rangle.$ (121)
Let
$\mathcal{M}_{1}=|\psi_{c}\rangle\langle\Omega|$
and
$\mathcal{M}_{2}=|\Omega\rangle\langle\psi_{c}|,$
then we only need to compute $\langle\psi_{1}|e^{iHt}\mathcal{M}_{1}e^{-iHt}|\psi_{1}\rangle$ and $\langle\psi_{1}|e^{iHt}\mathcal{M}_{2}e^{-iHt}|\psi_{1}\rangle$ to simulate the quantum Hadamard Test algorithm. Taking $\langle\psi_{1}|e^{iHt}\mathcal{M}_{1}e^{-iHt}|\psi_{1}\rangle$ as an example, noting that $|\psi_{c}\rangle=\sum_{\bm{j}}a_{\bm{j}}|\bm{j}\rangle$ represents a classical state with $R$ configurations where each configuration (product state) $|\bm{j}\rangle$ has $P$ particles, and amplitude $\left|a_{\bm{j}}\right|\geq 1/{\rm poly}(n)$. Then the operator
$\mathcal{M}_{1}=\sum_{\bm{j}}a_{\bm{j}}|\bm{j}\rangle\langle\Omega|,$ (122)
and we have
$\langle\psi_{1}|e^{iHt}\mathcal{M}_{1}e^{-iHt}|\psi_{1}\rangle=\sum_{\bm{j}}a_{\bm{j}}\langle\psi_{1}|e^{-Ht}|\bm{j}\rangle\langle\Omega|e^{-iHt}|\psi_{1}\rangle.$ (123)
Here, Alg. 1 has the ability to provide an $\epsilon/R$-approximation to each term $\langle\psi_{1}|e^{-Ht}|\bm{j}\rangle\langle\Omega|e^{-iHt}|\psi_{1}\rangle$, as a result, an $\mathcal{O}(\epsilon)$-approximation to the mean value $\langle\mathcal{M}_{1}\rangle$ is obtained with
$\mathcal{O}\left(R\left(\frac{2Rn}{\epsilon}\right)^{e^{2\pi te\mathfrak{d}}\log(2Rn/\epsilon)+\mathcal{O}(1)}\right)$ (124)
classical running time, where $R$ represents the number of involved configurations in the classical initial state $|\psi_{c}\rangle$.
Here, we note that the proposed algorithm does not necessarily limit to Hermitian observables. Without loss of generality, let $\mathcal{M}_{1,\bm{j}}=|1^{P}\rangle\langle 1^{P}|\otimes|0^{n-P}\rangle\langle 0^{n-P}|$ and $\mathcal{M}_{1,\bm{j}}^{(i)}\in\{|1\rangle\langle 0|,|0\rangle\langle 0|\}$. Lemma 4 can provide an estimation to $V_{i}(t)=e^{iHt}\mathcal{M}_{1,\bm{j}}^{(i)}e^{-iHt}$ by using the cluster expansion method, given by Eq. 38 which only requires the property local operator $\|\mathcal{M}_{1,\bm{j}}^{(i)}\|\leq 1$, but not restricted to the Hermitian observable. After obtaining local approximations $\{V_{1}(t),\cdots,V_{n}(t)\}$, we need to divide these operators into two regions $R_{1}$ and $R_{2}$, then compute $\langle\psi_{1}|V_{1}(t)\cdots V_{n}(t)|\psi_{1}\rangle=\langle\psi_{1}|V(R_{1})V(R_{2})|\psi_{1}\rangle$ according to Eq. 6. Combined with the quantum algorithm and Theorem 2 proposed by Ref [48], we conclude the result as follows.
Corollary 6 (Formal version of Corollary 2).
Given a 2D geometry local Hamiltonian satisfies certain symmetry, and a corresponding classical initial state $|\psi_{c}\rangle$ with $R$ configurations which has $p_{0}$ overlap to the ground state. Then there exists a classical algorithm that can output $\delta$-approximation to the ground state energy with the run time of
$\mathcal{O}\left(R\left(2Rn\right)^{e^{2\pi e\mathfrak{d}f(p_{0},\delta)}\log(2Rn)+\mathcal{O}(1)}\right),$ (125)
where $f(p_{0},\delta)\leq\mathcal{O}(\delta^{-1}\log(\delta^{-1}p_{0}^{-1}))$ and $\mathfrak{d}$ represents the maximum degree of $H$.
L.2 Eigenvalue Estimation
Given the Hamiltonian $H$ with certain symmetry property, we assume the classical initial state $|\psi_{c}\rangle$ contains several dominant modes. Specifically, let $\{(\lambda_{m},|\phi_{m}\rangle)\}_{m=1}^{2^{n}}$ represent pairs of eigenvalues and eigenvectors of $H$. We define $p_{m}=\left|\langle\phi_{m}|\psi_{c}\rangle\right|^{2}$ as the overlap between the initial state and the $m$-th eigenvector. Here, we follow the “Sufficiently Dominant Condition” assumption used in Ref. [78]: there exists a set of indices $\mathcal{D}\subset[2^{n}]$ such that $p_{\min}=\min_{i\in\mathcal{D}}p_{i}>p_{\rm tail}=\sum_{i\in\mathcal{D}^{c}}p_{i}$, where $\mathcal{D}^{c}=\{1,2,\cdots,M\}\setminus\mathcal{D}$.
Here, we follow the fundamental algorithms steps given by Ref. [78], but substitute the classical simulation algorithm into the quantum Hadamard test quantum circuit. In detail, the algorithm starts from (1) generating a proper set of $\{t_{k}\}_{k=1}^{N}$ according to some truncated Gaussian density; (2) Execute the classical simulation algorithm to simulate $Z_{k}=\langle\psi_{c}|e^{-iHt_{k}}|\psi_{c}\rangle$ with time $t_{k}$, and obtain the dataset $\{t_{k},Z_{k}\}$; (3) Classically post process $Z_{k}$ to derive the estimation for dominant eigenvalues $\{\lambda_{m}\}_{m\in\mathcal{D}}$. Combine the complexity result given by Eq. 124 and Theorem 3.1 in Ref. [78], we conclude the following result.
Corollary 7.
Given a $2$D geometry local Hamiltonian $H$ with eigenvalues and eigenvectors $\{\lambda_{m},|\psi_{m}\rangle\}$, and a classical initial state $|\psi_{c}\rangle$ enabling $p_{\min}>p_{\rm tail}$, there exists a classical algorithm that provides $\delta$-estimations to dominant eigenvalues within $|\psi_{c}\rangle$ such that
$\left|\lambda_{m}-\hat{\lambda}_{m}\right|\leq\delta$ (126)
for $m\in\mathcal{D}$. In particular, given the failure probability $\eta>0$, the classical algorithm takes
$\mathcal{O}\left(NR\left(2Rn\right)^{e^{2\pi e\mathfrak{d}/\delta}\log(2Rn)+\mathcal{O}(1)}+\mathcal{O}(1)\left|\mathcal{D}\right|\right)$ (127)
classical running time, where parameters $R$ represents the configurations given by $|\psi_{c}\rangle$, $\epsilon=\alpha/T$, $\alpha=\Omega\left(\log^{1/2}((p_{\min}-p_{\rm tail})^{-1})\right)$, $N=\Omega\left(\frac{1}{(p_{\min}-p_{\rm tail})^{2}}\log((Tb^{-1}+\left|\mathcal{D}\right|)/\eta)\right)$, and $b=\mathcal{O}(\log^{1/2}(\frac{p_{\min}}{p_{\rm tail}+(p_{\min}-p_{\rm tail})/2}))$.
Above result implies any $2$D geometry local Hamiltonian problem can be solved by a quasi-polynomial classical algorithm when the accuracy $\epsilon\in\mathcal{O}(1)$ and a pretty good classical initial state satisfies the sufficiently dominant condition.
Appendix M Simulating Adiabatic Dynamics
In this section, we provide the proof of Theorem 3. Our simulation strategy for the constant-time adiabatic dynamics consists essentially of approximating the expectation value estimator through the time-dependent cluster expansion as given in Sec. B.5. Practically, two types of error appear in our approximation: taking a finite $M$ and truncating the summation of $m$ to some finite threshold $T$. For bounding the error in the expectation value, we can equivalently bound the error in the operator norm such that we require the truncated approximation of $O(t)$ with finite $M$ and $T$ is $\epsilon$-close to Eq. (28). For convenience, we take the abbreviation of Eq. (28) as $O(t)=\lim_{M\rightarrow\infty}\sum_{m=0}^{+\infty}\frac{(-it)^{m}}{m!M^{m}}F_{m}$. We then apply the triangle inequality for accounting for the two sources of error individually:
$\left\|O(t)-\sum_{m=0}^{T}\frac{(-it)^{m}}{m!M^{m}}F_{m}\right\|\leq\left\|O(t)-\lim_{M\rightarrow\infty}\sum_{m=0}^{T}\frac{(-it)^{m}}{m!M^{m}}F_{m}\right\|+\left\|\lim_{M\rightarrow\infty}\sum_{m=0}^{T}\frac{(-it)^{m}}{m!M^{m}}F_{m}-\sum_{m=0}^{T}\frac{(-it)^{m}}{m!\tilde{M}^{m}}F_{m}\right\|=:\epsilon_{1}+\epsilon_{2},$ (128)
[…]
Without loss of generality, we assume that the adiabatic process is slowly changing such that $\lambda^{\prime}_{X}(t)\leq 1,\forall X\in S$. We next discuss the value of $f(n_{m},t,X)$ and $\tilde{\lambda}_{X}(t)$. For this problem, we take a more specific form of $\lambda_{X}(t)$ in the adiabatic Hamiltonian such that
$\lambda_{X}(t)=\begin{cases}t\lambda_{X},&\text{ if }\lambda_{X}\in H_{0},\\ (1-t)\lambda_{X},&\text{ if }\lambda_{X}\in H_{1}.\end{cases}$ (129)
It is then straightforward to see that the derivative of $\lambda_{X}(t)$ gives
$\mathrm{d}\lambda_{X}(t)=\begin{cases}\lambda_{X}\mathrm{d}t,&\text{ if }\lambda_{X}\in H_{0},\\ -\lambda_{X}\mathrm{d}t,&\text{ if }\lambda_{X}\in H_{1}.\end{cases}$ (130)
As a result, we have
$|f(n_{m},t,X)|=\left|\frac{\partial z_{X}\left(\frac{n_{m}t}{M}\right)}{\partial z_{x}(t)}\right|=\frac{\left|\frac{n_{m}}{M}\mathrm{d}t\lambda_{X}\left(\frac{n_{m}}{M}t\right)+\frac{n_{m}}{M}t\mathrm{d}\lambda_{X}\left(\frac{n_{m}t}{M}\right)\right|}{\left|\mathrm{d}t\lambda_{X}(t)+t\mathrm{d}\lambda_{X}(t)\right|}.$ (131)
First, consider the case $\lambda_{X}\in H_{0}$, we find that
$|f(n_{m},t,X)|=\frac{\left|\frac{n_{m}}{M}\mathrm{d}t\lambda_{x}\left(\frac{n_{m}}{M}t\right)+\frac{n_{m}}{M}+\mathrm{d}\lambda_{X}\left(\frac{n_{n}t}{M}\right)\right|}{\left|t\lambda_{x}dt+t\lambda_{X}\mathrm{d}t\right|}=\left(\frac{n_{M}}{M}\right)^{2}\leq\left(\frac{M-1}{M}\right)^{2}.$ (132)
Also, $\tilde{\lambda}_{X}(t)$ gives
$|\tilde{\lambda}_{X}(t)|=|t\lambda_{X}+t\lambda_{X}|\leq 2t.$ (133)
Next, for $\lambda_{X}\in H_{1}$, we have
$|f(n_{m},t,X)|=\frac{\frac{n_{m}}{M}\mathrm{d}t\lambda_{X}\left(\frac{n_{m}}{M}t\right)+\frac{n_{m}}{M}t\mathrm{d}\lambda_{X}\left(\frac{n_{m}}{M}t\right)}{\left|\mathrm{d}t(1-t)\lambda_{X}-t\lambda_{X}\mathrm{d}t\right|}=\left|\frac{n_{m}}{M}\cdot\frac{1-2\frac{n_{m}}{M}t}{1-2t}\right|.$ (134)
As the total simulation time for the adiabatic process is $t_{\text{total}}=1$ (and also for other time $t$ that is sufficiently away from $\frac{1}{2}$), we can further bound $\lim_{t\rightarrow 1}\left|\frac{1-\frac{2n_{m}}{M}t}{1-2t}\right|=\left|1-\frac{2n_{m}}{M}\right|\leqslant 1$. Therefore, we arrive at $|f(n_{m},t,X)|\leq\frac{n_{m}}{M}\leq\frac{M-1}{M}$. Alongside, $\tilde{\lambda}_{X}(t)$ in such cases accordingly provide
$|\tilde{\lambda}_{X}(t)|=|(1-t)\lambda_{X}-t\lambda_{X}|\leq 2t-1.$ (135)
Eventually, we combine the two cases and conclude that
$\begin{cases}&|f(n_{m},t,X)|\leq\frac{M-1}{M};\\ &|\tilde{\lambda}_{X}(t)|\leq 2t.\end{cases}$ (136)
For $\epsilon_{1}$, let us consider the value of $F_{m}$, where we first take the relaxation that $|f(n_{m},t,X)|\leq 1$ from Eq. (136). For $F_{m}$, the nested commutator will result in at most $2^{m}$ terms, each of which has value no more than $(2t)^{m}$; the summation over $n_{m},\cdots,n_{1}$ involves $M^{m}$ terms; and according to [Lemma 1, Ref. [35]] the total number of clusters of size $m$ is bounded by $(e\mathfrak{d})^{m}$, where $\mathfrak{d}$ is the maximum degree of the interaction graph. Therefore, we have $\|F_{m}\|\leq(4e\mathfrak{d}Mt)^{m}\|O\|$. The error $\epsilon_{1}$ can then be bounded as
$\left\|\lim_{M\rightarrow\infty}\sum_{m=T+1}^{+\infty}\frac{(-it)^{m}}{m!M^{m}}F_{m}\right\|\leq\lim_{M\rightarrow\infty}\sum_{m=T+1}^{+\infty}\frac{t^{m}}{m!M^{m}}\|F_{m}\|\leq\sum_{m=T+1}^{+\infty}\frac{(4e\mathfrak{d}t^{2})^{m}}{m!}\|O\|,$ (137)
which converges to $\|O\|\frac{\left(|t|/t^{*}\right)^{2T+2}}{1-(|t|/t^{*})^{2}}$ when $|t|\leq t^{*}:=\frac{1}{2\sqrt{e\mathfrak{d}}}$. Hence, taking
$\begin{split}\|O\|\frac{\left(|t|/t^{*}\right)^{2T+2}}{1-(|t|/t^{*})^{2}}\leq\epsilon/2\end{split}$ (138)
gives us the desirable $T$:
$T\geq\frac{\log\left(\frac{\epsilon(t/t^{*}-1)}{2\|O\|}\right)}{2\log(t^{*}/t)}-1.$ (139)
For the analysis of $\epsilon_{2}$, consider the truncation order of $M$ to be $\tilde{M}$. We find that
$\left\|\lim_{M\rightarrow\infty}\sum_{m=0}^{T}\frac{(-it)^{m}}{m!M^{m}}F_{m}-\sum_{m=0}^{T}\frac{(-it)^{m}}{m!\tilde{M}^{m}}F_{m}\right\|\leq\left|e^{4e\mathfrak{d}t^{2}}-e^{4e\mathfrak{d}t^{2}\left(\frac{\tilde{M}-1}{\tilde{M}}\right)}\right|\cdot\|O\|,$ (140)
then,
$\begin{split}\left|e^{4e\mathfrak{d}t^{2}}-e^{4e\mathfrak{d}t^{2}\left(\frac{\tilde{M}-1}{\tilde{M}}\right)}\right|\cdot\|O\|\leq\epsilon/2\\ \left|1-e^{-\frac{4e\mathfrak{d}t^{2}}{\tilde{M}}}\right|\leq e^{-4e\mathfrak{d}t^{2}}\frac{\epsilon}{2\|O\|}\\ \frac{4e\mathfrak{d}t^{2}}{\tilde{M}}\leq e^{-4e\mathfrak{d}t^{2}}\frac{\epsilon}{2\|O\|},\end{split}$ (141)
for which we reach
$\tilde{M}\geq\frac{8e\mathfrak{d}t^{2}\|O\|e^{4e\mathfrak{d}t^{2}}}{\epsilon}.$ (142)
Finally, we note that the complexity given Eq. (11) is delivered by the dominant term in the expansion, i.e., $\tilde{M}^{T}$.
This completes the proof.